### v2

# A theoretical analysis on the resolution of parametric PDEs via Neural Networks designed with Strassen algorithm

Abstract:
We construct a Neural Network that approximates the matrix multiplication operator for any activation function such that there exists a Neural Network which can approximate the scalar multiplication function. In particular, we use the Strassen algorithm to reduce the number of weights and layers needed for such Neural Networks. This allows us to define another Neural Network for approximating the inverse matrix operator. Also, by relying on the Galerkin method, we apply those Neural Networks to solve parametric elliptic PDEs for a whole set of parameters. Finally, we discuss improvements with respect to the prior results.
Key words: Approximation theory, Elliptic PDEs, Neural Networks, Parametric PDEs, Strassen algorithm.
Abbreviated title: Neural Networks and Strassen algorithm.
AMS subject classification: 35A35, 35J99, 41A25, 41A46, 65N30, 68T07
Acknowledgement: J.A.B.P. was supported by the Grant PID2021-126813NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. Also, both authors were supported by the grant IT1615-22 funded by the Basque Government.
1 Introduction
The Machine Learning explosion has left few research branches untouched, including Numerical Analysis for PDEs. An example are Physics Informed Neural Networks, which provide numerical solutions of PDEs (see [RPK19]). However, this is not the only example: [KPRS22] also constitutes an important breakthrough. There, the authors construct two $\operatorname{ReLU}$ Neural Networks capable of approximating the matrix multiplication and matrix inversion operators while giving an upper bound on the number of weights and layers (see Section 6 for the definition of $\operatorname{ReLU}$ activation function). These Neural Networks are then used in the Galerkin method for solving elliptic PDEs.
The cost of using these methods over conventional ones is hard to justify for solving a single particular PDE. However, they are useful to solve numerically parametric PDEs for a set of parameters. In fact, the use of Machine Learning for parametric PDEs tries to avoid the problems of solving them for each parameter. These problems are essentially the study of too many parameters or a high computing complexity for even a single PDE which both leads to an unreasonable time cost. When solving for a set of parameters at the same time, those problems are mitigated.
In this paper, we focus on parametric elliptic PDEs, which appear in many contexts such as physics and engineering. The problem of describing a physical phenomenon sometimes translates into solving an elliptic PDE. Many examples of this exist, such as the stationary Navier-Stokes, Poisson, Laplace, Helmholtz, and anisotropic heat diffusion equations. These equations explain fluid mechanics, electromagnetism, gravitational fields, and steady-state processes of heat/mass transfers respectively, as shown, for example, in [Pol01, PZ03]. Given their frequency and importance, it is necessary to understand the possible numerical algorithms to efficiently solve these equations.
1.1 Our Contribution
Motivated by the paper [KPRS22], we seek to understand how Neural Networks can approximate the solutions of parametric PDEs while improving the bounds on the number of parameters via mathematical arguments. For that, we also rely on the Galerkin method combined with the inversion of matrices obtained as the sum of a power series, though in a more precise way. Our contributions are the following ones:
  • Activation functions. In Neural Networks, the activation function plays a major role, as it determines its expressivity. In this paper, we focus on activation functions for which there are Neural Networks that approximate the product operation between two real numbers, as mathematically stated in Hypothesis 1. Clearly, if we cannot approximate such operation, there is no hope of representing the product of matrices, which is needed to apply afterwards the Galerkin method. This set of maps contains $\operatorname{ReLU}$, the activation function considered in [KPRS22].
  • Bounds. The authors of [KPRS22] bounded the number of parameters needed for Neural Networks used for parametric PDEs. To achieve more precise bounds, we use two key ideas. First, we use Strassen algorithm for matrix multiplication, which reduces the number of scalar multiplications needed for matrix multiplication. This allows us to economize the use of the potentially bulky Neural Network that approximates the product of scalars. Secondly, we use a finite sum of powers of matrices to approximate the inverses of matrices appearing in the Galerkin method. Rewriting that sum, we reduce the number of matrix multiplications in comparison to [KPRS22].
    In particular, under the hypothesis on the activation function, we prove two main theorems, which are compared to those in [KPRS22] in Section 6.
    • Theorem 3.4 provides a Neural Network that approximates the matrix multiplication operator. We achieve a smaller size of Neural Network than the one in [KPRS22] with a dependency on the number of rows of the input matrices that is no longer cubic, but of the order of $\log_{2}7$.
    • Theorem 4.6 provides a Neural Network that approximates the matrix inverse operator. Again, we achieve a smaller size of Neural Network than the one in [KPRS22] with a dependency on the number of rows of the input matrices that is no longer cubic, but of the order of $\log_{2}7$.
    As in [KPRS22], in Section 5, we see how Theorem 4.6 is used to numerically solve PDEs with the Galerkin method. The most common application of the Galerkin method, the finite element method, suffers from the curse of dimensionality, which we diminish with the exponent $\log_{2}7$.
  • Notation. Finally, we simplify notation by allowing matrices as input of Neural Networks. This avoids flattening matrices, which overloads the notation.
1.2 State of the art
In this subsection, we discuss the results of the literature related to this paper. The most related work is [KPRS22] since our work is a generalization and improvement of its results, as we have explained before.
For the rest of the literature review, in Section 1.2.1 we explore the matrix multiplication algorithms which optimize the amount of scalar multiplications. We follow in Section 1.2.2 by presenting the approximation theory for Neural Networks, which justifies that the product of scalars can be approximated by Neural Networks with a great variety of activation functions. Then, in Section 1.2.3 we explore the different interactions involving matrix multiplication, Neural Networks and the Strassen algorithm. Finally, in Section 1.2.4 we present the uses that Neural Networks receive to numerically solve PDEs.
1.2.1 Strassen-Like Algorithms
In [Str69], it is shown that for matrices of size $n\times n$, only $\mathcal{O}(n^{\log_{2}7})$ scalar multiplications are required, instead of $\theta(n^{3})$, which is what the definition suggests.
It is an open problem to find the infimum of $\omega$ such that for every $\varepsilon>0$ there is a $(n\times n)$-matrix multiplication algorithm with $\mathcal{O}(n^{\omega+\varepsilon})$ arithmetic operations. For that purpose a number of more optimal algorithms than Strassen appeared. We would like to highlight two of them: [CW90], which describes an algorithm that was not outperformed for 20 years, and [ADW+25], which is the latest and lowest upper bound. Both use the laser method introduced by Strassen in [Str87] that relates trilinear polynomials to the matrix multiplication tensor. This method, however, has a limitation: [AFLG15] shows that using that method and a variety of its generalizations cannot obtain the bound for $\omega=2$, which is a reasonable target.
Even if those papers present better algorithms than the one in [Str69], as the structure of the product is much more complicated or unknown, we leave it for future work.
1.2.2 Approximation Theory for Neural Networks
Approximation theory is critical to justify the success of Neural Networks. It studies its density and convergence to the desired function with the objective of answering when a Neural Network is suited for a given job.
The first results in approximation theory for Neural Networks are density theorems over $C^{0}$ and $L^{p}$ spaces, results known as Universal Approximation Theorems. The seminal work is [Cyb89], which proved the density of one layer Neural Networks with sigmoidal activation functions in the space of continuous functions equipped with the supremum norm. Generalizations followed, such as [HSW89], where a more general set of maps containing Neural Networks is proved to be dense in $C^{0}$, and such as [LLPS93], where the authors characterize for what activation functions the set of Neural Networks is dense in $L^{p}$ with $p\in(1,\infty)$. Density results are not limited to classic Lebesgue spaces. In fact, the author of [Hor91] gives conditions on the activation function ($m$ times continuously derivable and bounded) for the density in Sobolev spaces over compact and finite measures. As for Besov spaces, its relations with the approximation classes of Neural Network is reflected in [GKNV22]. However, in these papers, there is no computation on the number of weights required to approximate a given function.
In order to give upper bounds for the width and depth of Neural Network, some smoothness of the function is usually required. Examples of this are given in the papers [Bar93, Mha96, Yar17, PV18, Han19, GKP20, OPS20, GR21, Sie23]. The pioneer is [Bar93], which gives a relation between the number of weights of a one layer Neural Network and its proximity in $L^{2}$ norm to functions with an integrability condition on its Fourier Transform. For [Mha96], [Yar17], and [Sie23], the $L^{p}$ proximity of functions in $W^{s,p}$ with Neural Networks is expressed in terms of the number of weights and layers, and the Sobolev norm. In [OPS20], the authors do the same but for the Besov spaces. For $L^{p}$ norms, the authors of [PV18] shows the proximity of continuously differentiable to Neural Networks while tracking the necessary number of parameters. Moreover, given a continuous function, in [EHJ17] it is shown that there exists a sequence of Neural Networks with known number of parameters converging in the supreme norm to that function. In addition, in [GKP20] and [GR21], the authors provide upper bounds on the number of weights and layers to approximate in the Sobolev space $W^{s,p}$ by Neural Networks a target function which belongs to $W^{n,p}$, for $n>s$. Finally, the survey [GRK22] provides a general overview of this topic.
From a more computational perspective, the paper [DLL+19] proves that the most widely used algorithm for training Neural Networks, gradient descent, achieves the desired result as long as the activation function is smooth and Lipschitz. On the same topic, [AZLS19] guarantees, for the same algorithm and for sufficiently large Neural Networks, convergence, as well as providing estimates on its time cost.
1.2.3 Neural Networks and Multiplication
For scalar multiplication, in [Yar17] and [PV18] a $\operatorname{ReLU}$ Neural Network capable of approximating the product of any pair of scalars in an interval is described. Then, using that Neural Network as building block, the authors of [KPRS22] construct Neural Networks with $\operatorname{ReLU}$ as activation function which approximates the matrix multiplication and inverse operators. This is, as far as we know, the first explicitly constructed $\operatorname{ReLU}$ Neural Network to approximate matrix multiplication or its inverse with upper bounds on the number of weights and layers. For other activation functions like $\varrho(x)=x^{2}$ and $\operatorname{ReLU}^{2}$, those results are obvious while others, even if it is known to exist, there are no upper bounds on the number of parameters for a Neural Network that approximates scalar multiplication operator.
On the reciprocal problem of this work, Strassen-type matrix multiplication algorithms have been obtained using Neural Networks, by [Els16, FBH+22] among others. Additional examples of those algorithms can be found in the review
[MKA24]. However, it should be noted that the algorithms mentioned in Section 1.2.1 are more efficient asymptotically.
Optimization of evaluations and training of Neural Networks and other related architectures of Machine Learning is a huge and ever-evolving research topic. Among many other algorithms, the Strassen-type matrix multiplication algorithms are used in this context. This has been done from a purely computational perspective, for example, in [RR18, ZWWL18, AYK+20]. We are unaware of theoretical analysis of this kind of algorithms involving Strassen-type matrix multiplications.
1.2.4 Neural Networks and PDEs
Machine Learning techniques for numerically solving PDEs are a growing research field. The most relevant of this trend is the Physics Informed Neural Networks introduced in [SS18] and [RPK19]. A Physics Informed Neural Network is a Neural Network that approximates the solution of a PDE by minimizing its value to satisfy the equations on a sample of points in the domain. Other examples of using Neural Networks that are not Physics Informed Neural Networks to approximate numerically solutions of specific PDEs are in [EHJ17, HJKN20].
More recently, the aim of architectures has shifted from approximating particular solutions of a PDE to approximating its solver operator. One of the first of this kind to appear is the Deep Operator Net in [LJK19]. In fact, Neural Operators are a promising new approach for this task as explained in [KLL+23].
There is also a clash between the classical algorithm for numerically solving PDEs with this new wave of Machine Learning based algorithm. This results in mixing the two approaches to mitigate the computational costs of both techniques. Such classic algorithms are the Galerkin method and the Finite Element Method. Combining those methods, which requires an appropriate selection of functions, with Neural Networks, we get results such as those in [HLXZ20, LC20, DL21, ZCL+21, GPR+21]. The idea is to replace the hard-to-find functions that form a basis by a trainable family of Neural Networks. Other ideas are exploited in [KPRS22] with the aim of mathematical understanding instead of computational efficiency. Finally, mixing the two approaches is not limited either to Neural Networks. For example, the authors of [GZW22] propose Graph Convolutional Networks for this task.
1.3 Outline
The outline of the paper is the following: in Section 2, we fix the notation, describe the Strassen Algorithm, and define Matrix Neural Networks and related concepts. In Section 3, we construct Matrix Neural Networks that approximate the matrix multiplication operator. In Section 4, we construct a Matrix Neural Network that approximates the inverse of a set of matrices. In Section 5, we explain the use of the previous Matrix Neural Network for numerically solving PDEs. Finally, in Section 6, we compare our results with those in [KPRS22].
2 Preliminaries
2.1 Linear Algebra
In this paper, we denote the spaces of matrices of $n_{1}$ rows and $n_{2}$ columns by $\mathbb{R}^{n_{1}\times n_{2}}$. Given $A\in\mathbb{R}^{n_{1}\times n_{2}}$, $i\in\left\{1,\dots,n_{1}\right\}$ and $j\in\left\{1,\dots,n_{2}\right\}$, $A_{i,j}$ denotes the entry in its $i$-th row and $j$-th column. Moreover, given $c_{1},c_{2}\in\left\{1,\dots,n_{1}\right\}$ such that $c_{1}\leq c_{2}$ and $c_{3},c_{4}\in\left\{1,\dots,n_{2}\right\}$ such that $c_{3}\leq c_{4}$ we denote by $A_{c_{1}:c_{2},c_{3}:c_{4}}$ its submatrix formed by the rows from $c_{1}$ to $c_{2}$ (including $c_{1}$ and $c_{2}$), and the columns from $c_{3}$ to $c_{4}$ (including $c_{3}$ and $c_{4}$). We also denote by $\operatorname{I}_{n}\in\mathbb{R}^{n\times n}$ the identity matrix whose entries $(\operatorname{I}_{n})_{i,j}=\delta_{i,j}$ for all $i,j\in\left\{1,\dots,n\right\}$.
While working with Neural Networks, it is important to formalize the notion of matrix concatenation which we define as follows:
Definition 2.1.
Given two matrices $A\in\mathbb{R}^{n_{1}\times n_{2}}$ and $B\in\mathbb{R}^{n_{1}\times n_{3}}$, we define by $A|B$ the matrix in $\mathbb{R}^{n_{1}\times(n_{2}+n_{3})}$ whose first $n_{2}$ columns form the matrix $A$, and whose last $n_{3}$ columns form the matrix $B$; that is,
$(A|B)_{i,j}=\begin{cases}A_{i,j},&j\leq n_{2},\\ B_{i,j-n_{2}},&j>n_{2}.\end{cases}$
Similarly, given $A\in\mathbb{R}^{n_{1}\times n_{3}}$ and $B\in\mathbb{R}^{n_{2}\times n_{3}}$, the matrix $\begin{pmatrix}A\\ B\end{pmatrix}$ denotes the matrix in $\mathbb{R}^{(n_{1}+n_{2})\times n_{3}}$ whose first $n_{1}$ rows form the matrix $A$, and whose last $n_{2}$ rows form the matrix $B$; that is,
$\begin{pmatrix}A\\ B\end{pmatrix}_{i,j}=\begin{cases}A_{i,j},&i\leq n_{1},\\ B_{i-n_{1},j},&i>n_{1}.\end{cases}$
In addition, we consider the following notation that is especially useful for describing the Strassen algorithm:
Definition 2.2.
Let $A\in\mathbb{R}^{2^{k+1}\times 2^{k+1}}$ for some $k\in\mathbb{N}\cup\left\{0\right\}$. Then, we denote:
$A[1,1]$ :-$\coloneq A_{1:2^{k},1:2^{k}},:$ $A[1,2]$ :-$\coloneq A_{1:2^{k},2^{k}+1:2^{k+1}},:$
$A[2,1]$ :-$\coloneq A_{2^{k}+1:2^{k+1},1:2^{k}},:$ $A[2,2]$ :-$\coloneq A_{2^{k}+1:2^{k+1},2^{k}+1:2^{k+1}}.:$
Note that we have:
$A=\begin{pmatrix}A[1,1]|A[1,2]\\ A[2,1]|A[2,2]\end{pmatrix}.$
Throughout this paper, we consider linear functionals between spaces of matrices. In particular, we consider tensors located in $\mathbb{R}^{(n_{1}\times n_{2})\times(n_{3}\times n_{4})}$. We recall that, given $\mathcal{L}\in\mathbb{R}^{(n_{1}\times n_{2})\times(n_{3}\times n_{4})}$ and $A\in\mathbb{R}^{n_{3}\times n_{4}}$, the product $\mathcal{L}A$ is the matrix in $\mathbb{R}^{n_{3}\times n_{4}}$ given by:
$(\mathcal{L}A)_{i,j}=\sum_{l=1}^{n_{4}}\sum_{k=1}^{n_{3}}\mathcal{L}_{i,j,k,l}A_{k,l}.$
In particular, when $n_{1}\times n_{2}=n_{3}\times n_{4}$, we define the identity tensor $\operatorname{id}$, whose entries are:
$(\operatorname{id})_{i,j}=\delta_{i\times j,k\times l}.$
An important notion when working with Neural Network is the notion of weights; that is, of non null terms:
Definition 2.3.
Let $\mathcal{L}\in\mathbb{R}^{(n_{1}\times n_{2})\times(n_{3}\times n_{4})}$. We denote the number of weights of $\mathcal{L}$ by:
$\operatorname{n_{W}}\left(\mathcal{L}\right)=\#\left\{(i,j,k,l)\in\left\{1,\dots,n_{1}\right\}\times\left\{1,\dots,n_{2}\right\}\times\left\{1,\dots,n_{3}\right\}\times\left\{1,\dots,n_{4}\right\}\right.\\ \left.\;\middle|\;\mathcal{L}_{i,j,k,l}\neq 0\right\}.$
Similarly, if $A\in\mathbb{R}^{n_{1}\times n_{2}}$, we denote the number of weights of $A$ by:
$\operatorname{n_{W}}\left(A\right)=\#\left\{(i,j)\in\left\{1,\dots,n_{1}\right\}\times\left\{1,\dots,n_{2}\right\}\;\middle|\;A_{i,j}\neq 0\right\}.$
Finally, if $b\in\mathbb{R}^{n}$, we denote the number of weights of $b$ by:
$\operatorname{n_{W}}\left(b\right)=\#\left\{i\in\left\{1,\dots,n\right\}\;\middle|\;b_{i}\neq 0\right\}.$
In this paper ##\## denotes the cardinality of a set.
We will also need the following norms:
Definition 2.4.
Let $p\in[1,\infty]$, $A\in\mathbb{R}^{n_{1}\times n_{2}}$ and $x\in\mathbb{R}^{n_{1}}$. Then, we denote the $p$-norm of $A$ and $x$ by:
$\left\lvert A\right\rvert_{p}\coloneq\begin{cases}\left(\sum\limits_{i=1}^{n_{1}}\sum\limits_{j=1}^{n_{2}}\left\lvert A_{i,j}\right\rvert^{p}\right)^{\frac{1}{p}}&\text{ if }p<\infty,\\ \sup\limits_{\begin{subarray}{c}i=1,\dots,n_{1}\\ j=1,\dots,n_{2}\end{subarray}}\left\lvert A_{i,j}\right\rvert&\text{ if }p=\infty,\end{cases}$ $\left\lvert x\right\rvert_{p}=\begin{cases}\left(\sum\limits_{k=1}^{n_{1}}\left\lvert x_{i}\right\rvert^{p}\right)^{\frac{1}{p}}&\text{ if }p<\infty,\\ \sup\limits_{i=1,\dots,n_{1}}\left\lvert x_{i}\right\rvert&\text{ if }p=\infty.\end{cases}$
It should not be confused with induced norms, even if norms are equivalent in finite-dimensional spaces:
Definition 2.5.
Let $A\in\mathbb{R}^{n_{1}\times n_{2}}$ be a matrix, $p\in[1,\infty]$ and consider the $p$-norm $\left\lvert\;\cdot\;\right\rvert_{p}$ defined as in Definition 2.4. Then, we denote the induced norm of $A$ by:
$\left\lVert A\right\rVert_{p}\coloneq\sup_{x\neq 0}\frac{\left\lvert Ax\right\rvert_{p}}{\left\lvert x\right\rvert_{p}}.$
We recall the following estimates between different norms:
Lemma 2.6.
Let $A\in\mathbb{R}^{n\times n}$. Then,
$\left\lvert A\right\rvert_{\infty}\leq\left\lVert A\right\rVert_{2}\leq n\left\lvert A\right\rvert_{\infty}.$ (2.1)
2.2 Strassen Algorithm for Square Matrices
In this section, we recall the Strassen algorithm, which was proposed in [Str69]. First, we define the Strassen algorithm in matrices belonging to $\mathbb{R}^{2^{k}\times 2^{k}}$. Then, we define it in matrices in $\mathbb{R}^{n\times n}$ for arbitrary values $n\in\mathbb{N}$.
Let us now recursively define the Strassen algorithm for the matrices in $\mathbb{R}^{2^{k}\times 2^{k}}$. The base case $k=0$ simply consists of multiplications of scalars. The inductive case, $k\geq 1$, is defined as follows:
$(AB)[1,1]$ $=P_{1}+P_{4}-P_{5}+P_{7},$ $(AB)[1,2]$ $=P_{3}+P_{5},$ (2.2)
$(AB)[2,1]$ $=P_{2}+P_{4},$ $(AB)[2,2]$ $=P_{1}-P_{2}+P_{3}+P_{6},$
for:
$\begin{split}P_{1}&=(A[1,1]+A[2,2])(B[1,1]+B[2,2]),\\ P_{2}&=(A[2,1]+A[2,2])B[1,1],\\ P_{3}&=A[1,1](B[1,2]-B[2,2]),\\ P_{4}&=A[2,2](B[2,1]-B[1,1]),\\ P_{5}&=(A[1,1]+A[1,2])B[2,2],\\ P_{6}&=(A[2,1]-A[1,1])(B[1,1]+B[1,2]),\\ P_{7}&=(A[1,2]-A[2,2])(B[2,1]+B[2,2]),\end{split}$ (2.3)
where all the multiplications in (2.3) consist of the Strassen multiplication in $\mathbb{R}^{2^{k-1}\times 2^{k-1}}$.
When $n\neq 2^{k}$, the Strassen algorithm can be adapted in different ways:
  1. 1.
    We extend the matrix from $n\times n$ to $2^{k}\times 2^{k}$, for $k=\left\lceil\log_{2}(n)\right\rceil$ as follows:
    $A\mapsto\begin{bmatrix}A&0\\ 0&0\end{bmatrix}.$
    It is then enough to apply the Strassen multiplication to this extended matrix and recover the convenient submatrix.
  2. 2.
    We consider a division of blocks of size $2^{k}\times 2^{k}$, perform the multiplication with such blocks, and do the corresponding additions. For example, if $n=6$ and $A,B\in\mathbb{R}^{6\times 6}$, we can express the product $AB$ as a sum of products of submatrices where we can apply the Strassen algorithm:
    $(AB)_{1:4,1:4}$ $=A_{1:4,1:4}B_{1:4,1:4}+\begin{pmatrix}A_{1:2,5:6}B_{5:6,1:2}&A_{1:2,5:6}B_{5:6,3:4}\\ A_{3:4,5:6}B_{5:6,1:2}&A_{3:4,5:6}B_{5:6,3:4}\end{pmatrix},$
    $(AB)_{1:2,5:6}$ $=A_{1:2,1:2}B_{1:2,5:6}+A_{1:2,3:4}B_{3:4,5:6}+A_{1:2,5:6}B_{5:6,5:6},$
    $(AB)_{3:4,5:6}$ $=A_{3:4,1:2}B_{1:2,5:6}+A_{3:4,3:4}B_{3:4,5:6}+A_{3:4,5:6}B_{5:6,5:6},$
    $(AB)_{5:6,1:2}$ $=A_{5:6,1:2}B_{1:2,1:2}+A_{5:6,3:4}B_{3:4,1:2}+A_{5:6,5:6}B_{5:6,1:2},$
    $(AB)_{5:6,3:4}$ $=A_{5:6,1:2}B_{1:2,3:4}+A_{5:6,3:4}B_{3:4,3:4}+A_{5:6,5:6}B_{5:6,3:4},$
    $(AB)_{5:6,5:6}$ $=A_{5:6,1:2}B_{1:2,5:6}+A_{5:6,3:4}B_{3:4,5:6}+A_{5:6,5:6}B_{5:6,5:6}.$
2.3 Matrix Neural Networks
We now define the notions of Matrix Neural Network, concatenation, and parallelization. This is based on the definition introduced in [PV18], but adapted for Neural Networks whose entries are matrices. We do that for the sake of simplifying the notation.
Definition 2.7.
Let $L,N^{0,1},N^{0,2}\in\mathbb{N}$. A Matrix Neural Network $\Phi$ with input dimensions $\operatorname{dim_{in}}\Phi\coloneq N^{0,1}\times N^{0,2}$ and $L$ layers is a sequence of triplets of tensors, matrices and activation functions:
$\Phi=((\mathcal{L}^{1},C^{1},\alpha^{1}),(\mathcal{L}^{2},C^{2},\alpha^{2}),\dots,(\mathcal{L}^{L},C^{L},\alpha^{L}))$
where $N^{1,1},N^{1,2},\dots,N^{L,1},N^{L,2}\in\mathbb{N}$, and where each $\mathcal{L}^{\ell}$ is a
$(N^{\ell,1}\times N^{\ell,2})\times(N^{\ell-1,1}\times N^{\ell-1,2})$
tensor, $C^{\ell}$ is a $N^{\ell,1}\times N^{\ell,2}$ matrix and
[…]
for all $\ell=1,\dots,L$.
If $\Phi$ is a Matrix Neural Network as above and if $\varrho:\mathbb{R}\longrightarrow\mathbb{R}$, then we say that $\Phi$ is a $\varrho$-MNN if $\alpha^{\ell}_{i,j}\in\left\{\operatorname{id},\varrho\right\}$ for $i=1,\dots,N^{\ell,1}$, $j=1,\dots,N^{\ell,2}$, $\ell=1,\dots,L-1$, and $\alpha^{L}_{i,j}=\operatorname{id}$ for all $i=1,\dots,N^{L,1}$ and $j=1,\dots,N^{L,2}$.
Moreover, if $\Phi$ is a Matrix Neural Network as above, we define the realization of $\Phi$ as the map $R(\Phi):\mathbb{R}^{N^{0,1}\times N^{0,2}}\longrightarrow\mathbb{R}^{N^{L,1}\times N^{L,2}}$ such that
$R(\Phi)(A)=A^{L},$
where $A^{L}$ results from the following scheme:
$A^{0}$ :-$\coloneq A,:$
$A^{\ell}$ :-$\coloneq\alpha^{\ell}(\mathcal{L}^{\ell}A^{\ell-1}+C^{\ell}),\quad\text{ for }\ell=1,\dots,L.:$
We call $L=L(\Phi)$ the number of layers. For $\ell\leq K$ we call
$M_{\ell}(\Phi)\coloneq\operatorname{n_{W}}\left(\mathcal{L}^{\ell}\right)+\operatorname{n_{W}}\left(C^{\ell}\right)$
the number of weights in the $\ell$-th layer and we define
$M(\Phi)\coloneq\sum_{\ell=1}^{L}M^{\ell}(\Phi),$
which we call the number of weights of $\Phi$. Moreover, we refer to
$\operatorname{dim_{out}}\Phi\coloneq N^{L,1}\times N^{L,2}$
as the output dimension of $\Phi$.
Finally, given a previously defined Matrix Neural Network $\Phi$, in order to work with its layers, we write $((\mathcal{L}^{1},B^{1},\alpha^{1}),\dots,(\mathcal{L}^{L},B^{L},\alpha^{L}))\coloneq\Phi$ to specify which and in what order tensors, matrices, and activation functions conform to the Matrix Neural Network.
Remark 2.8.
Classical Neural Networks (if we consider them defined as in [PV18]) have the same structure as a Matrix Neural Networks: they are a sequence of triples of matrices, vectors, and activation functions. Their realization, number of layers, and number of weights are defined equivalently for these Neural Networks. In particular, any classical Neural Network can be viewed as a Matrix Neural Network by considering $N^{\ell,2}=1$ for all $\ell=1,\dots,N$. Conversely, using vectorization procedures, since the matrix space $\mathbb{R}^{n\times m}$ is isomorphic to the vector space $\mathbb{R}^{nm}$, all the Matrix Neural Networks can be implemented as classical Neural Networks (with the same length, number of weights, and final error). Therefore, we can view both notions as equivalent.
One important notion that we are going to use is the parallelization of Neural Networks:
Definition 2.9.
Let $L_{1},L_{2}\in\mathbb{N}$ and let
$\Phi^{1}$ $=\left((\mathcal{L}^{1,1},C^{1,1},\alpha^{1}),\dots,(\mathcal{L}^{1,L_{1}},C^{1,L_{1}},\alpha^{1,L_{1}})\right),$
$\Phi^{2}$ $=\left((\mathcal{L}^{2,1},C^{2,1},\alpha^{1}),\dots,(\mathcal{L}^{2,L_{2}},C^{2,L_{2}},\alpha^{2,L_{2}})\right)$
be two Matrix Neural Networks such that the input layer of $\Phi^{1}$ has the same dimension as the output layer of $\Phi^{2}$. Then, $\Phi^{1}\circ\Phi^{2}$ denotes the following $L_{1}+L_{2}$ layer Neural Network:
$\Phi^{1}\circ\Phi^{2}\coloneq\left((\mathcal{L}^{2,1},C^{2,1},\alpha^{2,1}),\dots,(\mathcal{L}^{2,L_{2}},C^{2,L_{2}},\alpha^{2,L_{2}}),\right.\\ \left.(\mathcal{L}^{1,1},C^{1,1},\alpha^{1,1}),\dots,(\mathcal{L}^{1,L_{1}},C^{1,L_{1}},\alpha^{1,L_{1}})\right)$
We call $\Phi^{1}\circ\Phi^{2}$ the concatenation of $\Phi^{1}$ and $\Phi^{2}$.
Remark 2.10.
In the literature, this concatenation is known as a sparse concatenation and is denoted by $\odot$. This difference comes from the fact that there is another kind of concatenation denoted by $\bullet$: given two $\varrho$-Matrix Neural Networks
$\Phi^{1}$ $=\left((\mathcal{L}^{1,1},C^{1,1},\alpha^{1}),\dots,(\mathcal{L}^{1,L_{1}},C^{1,L_{1}},\alpha^{1,L_{1}})\right),$
$\Phi^{2}$ $=\left((\mathcal{L}^{2,1},C^{2,1},\alpha^{1}),\dots,(\mathcal{L}^{2,L_{2}},C^{2,L_{2}},\alpha^{2,L_{2}})\right),$
such that $\operatorname{dim_{in}}\Phi^{1}=\operatorname{dim_{out}}\Phi^{2}$, the $\varrho$-Matrix Neural Network $\Phi^{1}\bullet\Phi^{2}$ is defined as:
$\Phi^{1}\bullet\Phi^{2}\coloneq\left((\mathcal{L}^{2,1},C^{2,1},\alpha^{2,1}),\dots,(\mathcal{L}^{2,L_{2}-1},C^{2,L_{2}-1},\alpha^{2,L_{2}-1}),\right.\\ \left.(\mathcal{L}^{1,1}\mathcal{L}^{2,L_{2}},\mathcal{L}^{1,1}C^{2,L_{2}}+C^{1,1},\alpha^{1,1}),\dots,(\mathcal{L}^{1,L_{1}},C^{1,L_{1}},\alpha^{1,L_{1}})\right).$
This $\varrho$-MNN satisfies again the property $R\left(\Phi^{1}\bullet\Phi^{2}\right)=R\left(\Phi^{1}\right)\circ R\left(\Phi^{2}\right)$. Even though this definition of concatenation produces Matrix Neural Networks with fewer layers, the estimation on the number of weights generally gets worse. This leads us to not use this definition.
Note that if $\Phi^{1}$ and $\Phi^{2}$ are a $\varrho$-MNN, then $\Phi^{1}\circ\Phi^{2}$ is also a $\varrho$-MNN too.
Definition 2.11.
Let $f^{i}:\mathbb{R}^{n_{i}\times n}\longrightarrow\mathbb{R}^{m_{i}\times m}$ for $i=1,2$ be two maps. Then, its tensor product $f^{1}\times f^{2}$ is the map defined as:
[…]
Remark 2.12.
It is easy to see that the tensor product of two tensors, viewed as maps between matrices spaces, is a tensor too.
Definition 2.13.
Let $L\in\mathbb{N}$ and let $\Phi^{1},\dots,\Phi^{k}$ be Neural Networks such that the input dimensions are shared and $\Phi^{i}=\left((\mathcal{L}^{i,1},C^{i,1},\alpha^{i,1}),\dots,(\mathcal{L}^{i,L},C^{i,L},\alpha^{i,L})\right)$. Then, we define the parallelization of $\Phi^{1},\dots,\Phi^{k}$ by:
$\mathcal{P}(\Phi^{1},\dots,\Phi^{k})\coloneq\left(\left(\mathcal{L}^{1,1}\times\dots\times\mathcal{L}^{k,1},\begin{pmatrix}C^{1,1}\\ \vdots\\ C^{k,1}\end{pmatrix},\alpha^{1,1}\times\dots\times\alpha^{k,1}\right),\dots,\right.\\ \left.\left(\mathcal{L}^{1,L}\times\dots\times\mathcal{L}^{k,L},\begin{pmatrix}C^{1,L}\\ \vdots\\ C^{k,L}\end{pmatrix},\alpha^{1,L}\times\dots\times\alpha^{k,L}\right)\right).$
In order to implement the Strassen algorithm with $\varrho$-MM, with $\varrho$ fixed, we consider that we may multiply two numbers using Neural Networks. In particular, we suppose the following.
Hypothesis 1.
Let $\varrho$ be an activation function. Then, for all $K,\varepsilon>0$ there is a $\varrho$-MNN $\Pi_{\varepsilon,K}^{\varrho}$ such that:
$\sup_{(x,y)\in[-K,K]^{2}}\left\lvert xy-R(\Pi_{\varepsilon,K}^{\varrho})(x,y)\right\rvert\leq\varepsilon.$
An important example is ReLU, as we show in Section 6.
Another relevant example is $\varrho(t)=\max\left\{0,t\right\}^{2}$. In this case, the product is a $\varrho$-realization:
$xy=\frac{1}{4}\left(\varrho(x+y)+\varrho(-x-y)-\varrho(x-y)-\varrho(-x+y)\right)$
We have used that:
$\varrho(t)+\varrho(-t)=t^{2}$
and
$(x+y)^{2}-(x-y)^{2}=4xy.$
Thus, for $\varrho(t)=\max\left\{0,t\right\}^{2}$ there is a Neural Network $\Pi^{\varrho}_{0,\infty}$ satisfying:
$M(\Pi^{\varrho}_{0,\infty})=12,\quad L(\Pi^{\varrho}_{0,\infty})=2$
and computes the product exactly. This network is:
$\Pi^{\varrho}_{0,\infty}=\left(\left(\begin{pmatrix}1&1\\ -1&-1\\ 1&-1\\ -1&1\end{pmatrix},0,\varrho\times\varrho\times\varrho\times\varrho\right),\left(\frac{1}{4}\begin{pmatrix}1&1&-1&-1\end{pmatrix},0,\operatorname{id}\right)\right).$
3 Strassen Neural Networks
3.1 Strassen Algorithm for Square Matrices Whose Size is a Power of 2 Using Neural Networks
We proceed to define the Strassen MNN to approximate the multiplication of matrices in $\mathbb{R}^{2^{k}\times 2^{k}}$ assuming Hypothesis 1. In particular, we consider Neural Networks that are defined recursively on $k$:
The base case is given by:
$R\left(\Phi^{\operatorname{STR}_{2^{0},\varrho}}_{\varepsilon,K}\right)(A|B)=R\left(\Pi_{\varepsilon,K}^{\varrho}\right)(A_{1,1},B_{1,1}),$ (3.1)
for $\Pi_{\varepsilon,K}^{\varrho}$ defined in Hypothesis 1.
For the recursive case, we need to define several auxiliary networks. The first, whose input is in $\mathbb{R}^{(7\cdot 2^{k-1})\times 2^{k-1}}$ and whose output is in $\mathbb{R}^{2^{k}\times 2^{k}}$, is given by:
$R\left(\Phi^{\operatorname{MIX}_{k}}\right)\begin{bmatrix}P_{1}\\ P_{2}\\ P_{3}\\ P_{4}\\ P_{5}\\ P_{6}\\ P_{7}\end{bmatrix}=\begin{pmatrix}P_{1}+P_{4}-P_{5}+P_{7}&P_{3}+P_{5}\\ P_{2}+P_{4}&P_{1}-P_{2}+P_{3}+P_{6}\end{pmatrix}.$
Note that $\Phi^{\operatorname{MIX}_{k}}$ is a one layer MNN such that:
$M\left(\Phi^{\operatorname{MIX}_{k}}\right)=12\cdot(2^{k-1})^{2}=3\cdot 4^{k}.$ (3.2)
Also, with the notation of Definition 2.2, we define a MNN whose input is in $\mathbb{R}^{2^{k}\times 2^{k+1}}$ and whose output is in $\mathbb{R}^{7\cdot 2^{k-1}\times 2\cdot 2^{k-1}}$:
$R(\Phi^{\operatorname{SPLIT}_{k}})(A|B)=\begin{pmatrix}A[1,1]+A[2,2]&|&B[1,1]+B[2,2]\\ A[2,1]+A[2,2]&|&B[1,1]\\ A[1,1]&|&B[1,2]-B[2,2]\\ A[2,2]&|&B[2,1]-B[1,1]\\ A[1,1]+A[1,2]&|&B[2,2]\\ A[2,1]-A[1,1]&|&B[1,1]+B[1,2]\\ A[1,2]-A[2,2]&|&B[2,1]+B[2,2]\end{pmatrix}.$
This network satisfies:
$M(\Phi^{\operatorname{SPLIT}_{k}})=24\cdot(2^{k-1})^{2}=6\cdot 4^{k}.$ (3.3)
Finally, we define a MNN whose input is in $\mathbb{R}^{7\cdot 2^{k-1}\times 2\cdot 2^{k-1}}$ and whose output is in $\mathbb{R}^{7\cdot 2^{k-1}\times 2^{k-1}}$:
$\Phi^{\operatorname{PAR}_{k,\varrho}}_{\varepsilon,K}=\mathcal{P}\left(\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\right.\\ \left.\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K}\right).$
Bearing this in mind, we define recursively that:
$\Phi^{\operatorname{STR}_{2^{k},\varrho}}_{\varepsilon,K}=\Phi^{\operatorname{MIX}_{k}}\circ\Phi^{\operatorname{PAR}_{k,\varrho}}_{\varepsilon,K}\circ\Phi^{\operatorname{SPLIT}_{k}}.$
Let us show the following result with respect to the accuracy and size of the MNN that we have designed:
Proposition 3.1.
Let $k\in\mathbb{N}\cup\left\{0\right\}$, $A,B\in\mathbb{R}^{2^{k}\times 2^{k}}$, $\varepsilon>0$, and $K>0$ satisfying:
$\left\lvert A\right\rvert_{\infty},\left\lvert B\right\rvert_{\infty}\leq K.$
Then,
$\left\lvert R\left(\Phi^{\operatorname{STR}_{2^{k},\varrho}}_{\varepsilon,K}\right)(A|B)-AB\right\rvert_{\infty}\leq\varepsilon,$ (3.4)
$M\left(\Phi^{\operatorname{STR}_{2^{k},\varrho}}_{\varepsilon,K}\right)=7^{k}\left(M\left(\Pi^{\varrho}_{\frac{\varepsilon}{4^{k}},2^{k}K}\right)+12\right)-12\cdot 4^{k},$ (3.5)
$L\left(\Phi^{\operatorname{STR}_{2^{k},\varrho}}_{\varepsilon,K}\right)=L\left(\Pi_{\frac{\varepsilon}{4^{k}},2^{k}K}^{\varrho}\right)+2k.$ (3.6)
Proof.
The statement (3.4) can be proved inductively on $k$. The base case $k=0$, when $A$ and $B$ are $1\times 1$ matrices, that is, scalars, is a direct consequence of (3.1) and Hypothesis 1.
Let us suppose that Proposition 3.1 is true for $k-1$. Let us start by showing (3.4). Using the Strassen algorithm, we know that (2.2) is satisfied. Consequently, because of the triangular inequality, defining $P_{i}$ as in (2.3), it suffices to show that:
[…]
Let us show the estimate for $i=1$, as the other six estimates are analogous. Since $\left\lvert A\right\rvert_{\infty}\leq K$, then $\left\lvert A[1,1]+A[2,2]\right\rvert_{\infty}\leq 2K$. Similarly, since $\left\lvert B\right\rvert_{\infty}\leq K$, we have $\left\lvert B[1,1]+B[2,2]\right\rvert_{\infty}\leq 2K$. Thus, using the recursive hypothesis with $k-1$ we obtain that:
[…]
concluding the proof of (3.4).
To continue with, let us show that the number of weights of the Matrix Neural Network is given by (3.5). We clearly have:
$M\left(\Phi_{\varepsilon,K}^{\operatorname{PAR}_{k,\varrho}}\right)=7M\left(\Phi_{\frac{\varepsilon}{4},2K}^{\operatorname{STR}_{2^{k-1},\varrho}}\right).$
Using the previous equality, (3.2) and (3.3), we obtain recursively:
$M\left(\Phi^{\operatorname{STR}_{2^{k},\varrho}}_{\varepsilon,K}\right)$ $=M\left(\Phi_{\varepsilon,K}^{\operatorname{PAR}_{k,\varrho}}\right)+M\left(\Phi^{\operatorname{MIX}_{k}}\right)+M\left(\Phi^{\operatorname{SPLIT}_{k}}\right)$ (3.8)
$=7M\left(\Phi_{\frac{\varepsilon}{4},2K}^{\operatorname{STR}_{2^{k-1},\varrho}}\right)+9\cdot 4^{k}$
$=7\left(7M\left(\Phi_{\frac{\varepsilon}{4^{2}},2^{2}K}^{\operatorname{STR}_{2^{k-2},\varrho}}\right)+9\cdot 4^{k-1}\right)+9\cdot 4^{k}$
=$\cdots$= ⋯
$=7^{k}M\left(\Pi^{\varrho}_{\frac{\varepsilon}{4^{k}},2^{k}K}\right)+9\sum_{i=1}^{k}4^{i}\cdot 7^{k-i}.$
We now calculate the sum using the fact that it is geometric:
$9\sum_{i=1}^{k}4^{i}\cdot 7^{k-i}=9\cdot 7^{k}\sum_{i=1}^{k}\left(\frac{4}{7}\right)^{i}=9\cdot 7^{k}\frac{4/7-(4/7)^{k+1}}{1-4/7}=12\left(7^{k}-4^{k}\right).$ (3.9)
Putting together (3.8) and (3.9), we obtain (3.5).
To finish, let us obtain the length of the Matrix Neural Network (3.6). For that purpose, we remark that for $k\geq 1$:
$L\left(\Phi^{\operatorname{STR}_{2^{k},\varrho}}_{\varepsilon,K}\right)=L\left(\Phi_{\frac{\varepsilon}{4},2K}^{\operatorname{STR}_{2^{k-1},\varrho}}\right)+2.$
Thus, using induction, we conclude with (3.6). ∎
3.2 Strassen Algorithm for Arbitrary Square Matrices Using Neural Networks
Consider $m,n,p\in\mathbb{N}$. In order to multiply a matrix $A\in\mathbb{R}^{m\times n}$ by another matrix $B\in\mathbb{R}^{n\times p}$ using the Strassen algorithm implemented with Neural Networks, we are going to prolong the matrices to size $2^{k}\times 2^{k}$, for:
$k=k_{m,p}^{n}\coloneq\left\lceil\log_{2}(\max\left\{m,n,p\right\})\right\rceil.$
In particular, we recall the estimate:
$\log_{2}(\max\left\{m,n,p\right\}))<k_{m,p}^{n}=\left\lceil\log_{2}(\max\left\{m,n,p\right\}))\right\rceil\leq\log_{2}(\max\left\{m,n,p\right\})+1.$ (3.10)
For that purpose, we define a Neural Network whose input is in $\mathbb{R}^{n\times(m+p)}$ and whose output is in $\mathbb{R}^{2^{k}\times 2\cdot 2^{k}}$ with $k=k_{m,p}^{n}$:
$R\left(\Phi^{\operatorname{EXT}_{n}}_{m,p}\right)\left(A^{T}\middle|B\right)=\left.\begin{pmatrix}A&0^{m,2^{k}-n}\\ 0^{2^{k}-m,n}&0^{2^{k}-m,2^{k}-n}\end{pmatrix}\middle|\begin{pmatrix}B&0^{n,2^{k}-p}\\ 0^{2^{k}-n,p}&0^{2^{k}-n,2^{k}-p}\end{pmatrix}\right.\in\mathbb{R}^{2^{k}\times 2\cdot 2^{k}},$
where $A^{T}$ denotes the transpose of $A$ and $0^{m,n}=(0)_{\begin{subarray}{c}i=1,\dots,m\\ j=1,\dots,n\end{subarray}}\in\mathbb{R}^{m\times n}$. Transposing the left-hand side matrix is necessary to obtain a well-defined matrix as input. Similarly, we define a Neural Network whose input is in $\mathbb{R}^{2^{k_{m,p}^{n}}\times 2^{k_{m,p}^{n}}}$ and output is in $\mathbb{R}^{m\times n}$:
$R\left(\Phi_{m,p}^{\operatorname{SHR}_{n}}\right)(A)=A_{1:m,1:p}\in\mathbb{R}^{m\times p}.$
It is easy to check that:
$M\left(\Phi^{\operatorname{EXT}_{n}}_{m,p}\right)=n(m+p)$ and $M\left(\Phi_{m,p}^{\operatorname{SHR}_{n}}\right)=mp.$ (3.11)
Thus, we can define the Neural Network:
$\Phi_{\varepsilon,K}^{\operatorname{STR}_{m,p,\varrho}^{n}}=\Phi^{\operatorname{SHR}_{n}}_{m,p}\circ\Phi_{\varepsilon,K}^{\operatorname{STR}_{2^{k_{m,p}^{n}},\varrho}}\circ\Phi^{\operatorname{EXT}_{n}}_{m,p}.$
Proposition 3.2.
Let $A\in\mathbb{R}^{m\times n}$, $B\in\mathbb{R}^{n\times p}$ and $K>0$ satisfying:
$\left\lvert A\right\rvert_{\infty},\left\lvert B\right\rvert_{\infty}\leq K.$
Then, for every $\varepsilon>0$,
$\left\lvert R\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{m,p,\varrho}^{n}}\right)\left(A^{T}\middle|B\right)-AB\right\rvert_{\infty}\leq\varepsilon,$ (3.12)
$M\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{m,p,\varrho}^{n}}\right)\leq 7\max\left\{m,n,p\right\}^{\log_{2}7}\left(M\left(\Pi^{\varrho}_{\frac{\varepsilon}{4\max\left\{m,n,p\right\}^{2}},2\max\left\{m,n,p\right\}K}\right)+12\right)$ (3.13)
$\phantom{aaaaaaaaaaaaa}-9\max\left\{m,n,p\right\}^{2},$
$L\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{m,p,\varrho}^{n}}\right)\leq L\left(\Pi_{\frac{\varepsilon}{4\max\left\{m,n,p\right\}^{2}},2\max\left\{m,n,p\right\}K}^{\varrho}\right)+2(\log_{2}(\max\left\{m,n,p\right\})+2).$ (3.14)
Proof.
Considering (3.4), we clearly have (3.12).
In addition, the number of weights can be obtained using (3.5), (3.11) and (3.10):
$M\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{m,p,\varrho}^{n}}\right)$ $=mp+M\left(\Phi^{\operatorname{STR}_{2^{k^{n}_{m,p}},\varrho}}_{\varepsilon,K}\right)+n(m+p)$
$=7^{k^{n}_{m,p}}\left(M\left(\Pi^{\varrho}_{\varepsilon/4^{k^{n}_{m,p}},2^{k^{n}_{m,p}}K}\right)+12\right)-12\cdot 4^{k^{n}_{m,p}}+mp+mn+np$
$\leq 7\max\left\{m,n,p\right\}^{\log_{2}7}\left(M\left(\Pi^{\varrho}_{\frac{\varepsilon}{4\max\left\{m,n,p\right\}^{2}},2\max\left\{m,n,p\right\}K}\right)+12\right)$
$-12\max\left\{m,n,p\right\}^{2}+mp+mn+np$
$\leq 7\max\left\{m,n,p\right\}^{\log_{2}7}\left(M\left(\Pi^{\varrho}_{\frac{\varepsilon}{4\max\left\{m,n,p\right\}^{2}},2\max\left\{m,n,p\right\}K}\right)+12\right)$
$-9\max\left\{m,n,p\right\}^{2}.$
To finish, regarding the length of the Neural Network, considering (3.6), (3.10), and that we have added two shallow Neural Networks:
$L\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{m,p,\varrho}^{n}}\right)$ $=L\left(\Phi_{m,p}^{\operatorname{SHR}_{n}}\right)+L\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{2^{k_{m,p}^{n}},\varrho}}\right)+L\left(\Phi_{m,p}^{\operatorname{EXT}_{n}}\right)$
$=L\left(\Pi_{\varepsilon/4^{k_{m,p}^{n}},2^{k_{m,p}^{n}}K}^{\varrho}\right)+2(k_{m,p}^{n}+1)$
$\leq L\left(\Pi_{\frac{\varepsilon}{4\max\left\{m,n,p\right\}^{2}},2\max\left\{m,n,p\right\}K}^{\varrho}\right)+2(\log_{2}(\max\left\{m,n,p\right\})+2).$
When we consider square matrices, we can extend the definition of the Neural Network $\Phi^{\operatorname{STR}_{n,\varrho}}_{\varepsilon,K}$ to $n\neq 2^{k}$ as follows:
$\Phi^{\operatorname{STR}_{n,\varrho}}_{\varepsilon,K}=\Phi_{n,n}^{\operatorname{SHR}_{n}}\circ\Phi_{\varepsilon,K}^{\operatorname{STR}_{2^{k},\varrho}}\circ\Phi_{n}^{\operatorname{EXT}^{*}},$ (3.15)
where $k=\left\lceil\log_{2}n\right\rceil$ and $\Phi_{n}^{\operatorname{EXT}^{*}}$ is a MNN whose input is in $\mathbb{R}^{n\times 2n}$ and output in $\mathbb{R}^{2^{k}\times 2\cdot 2^{k}}$ satisfying:
$R\left(\Phi_{n}^{\operatorname{EXT}^{*}}\right)(A|B)=\left.\begin{pmatrix}A&0^{n,n-2^{k}}\\ 0^{2^{k}-n,n}&0^{2^{k}-n,2^{k}-n}\end{pmatrix}\middle|\begin{pmatrix}B&0^{n,n-2^{k}}\\ 0^{2^{k}-n,n}&0^{2^{k}-n,2^{k}-n}\end{pmatrix}\right.$
for all $A,B\in\mathbb{R}^{n\times n}$. Note that we have the following equality:
$M\left(\Phi_{n}^{\operatorname{EXT}^{*}}\right)=2n^{2}.$
Remark 3.3.
When $n=2^{k}$, $R\left(\Phi^{SHR_{n}}_{n,n}\right)$ and $R\left(\Phi_{n}^{EXT^{*}}\right)$ are identity tensors, which is why we consider (3.15) a generalization. In that case, there is only a small difference in its length and number of weights, which is irrelevant for studying their asymptotic values.
Considering Proposition 3.2, we have that:
Theorem 3.4.
Let $A,B\in\mathbb{R}^{n\times n}$, and $K>0$ satisfying:
$\left\lvert A\right\rvert_{\infty},\left\lvert B\right\rvert_{\infty}\leq K.$
Then, for every $\varepsilon>0$,
$\left\lvert R\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{n,\varrho}}\right)\left(A\middle|B\right)-AB\right\rvert_{\infty}\leq\varepsilon,$ (3.16)
$M\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{n,\varrho}}\right)\leq 7n^{\log_{2}7}\left(M\left(\Pi^{\varrho}_{\frac{\varepsilon}{4n^{2}},2nK}\right)+12\right)-9n^{2},$ (3.17)
$L\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{n,\varrho}}\right)\leq L\left(\Pi_{\frac{\varepsilon}{4n^{2}},2nK}^{\varrho}\right)+2(\log_{2}n+2).$ (3.18)
Indeed, the resulting Matrix Neural Network is as in Proposition 3.2 but without transposing.
4 A Neural Network for Matrix Invertion
In this section, we aim to construct a Matrix Neural Network which, given a matrix input, returns an approximation to its inverse. This can be achieved using our newly defined Strassen Neural Networks to approximate the matrix product and the Neumann series for matrices.
We know that for all $A\in\mathbb{R}^{n\times n}$ such that $\left\lVert A\right\rVert_{2}<1$ (recall Definition 2.5), we have the identity:
$(\operatorname{I}_{n}-A)^{-1}=\sum_{k=0}^{\infty}A^{k},$
where the right-hand side is called the Neumann series of $A$. We also have the estimate:
$\left\lVert(\operatorname{I}_{n}-A)^{-1}-\sum_{k=0}^{N}A^{k}\right\rVert_{2}=\left\lVert\sum_{k=N+1}^{\infty}A^{k}\right\rVert_{2}\leq\sum_{k=N+1}^{\infty}\left\lVert A\right\rVert^{k}_{2}=\frac{\left\lVert A\right\rVert_{2}^{N+1}}{1-\left\lVert A\right\rVert_{2}}.$ (4.1)
Consequently, we can approximate the Neumann series of a matrix computing the partial sums while minimizing the number of matrix products using the following identity:
$\sum_{k=0}^{2^{N+1}-1}A^{k}=\left(A^{2^{N}}+\operatorname{I}_{n}\right)\sum_{k=0}^{2^{N}-1}A^{k}=\prod_{k=0}^{N}\left(A^{2^{k}}+\operatorname{I}_{n}\right).$ (4.2)
Indeed, the left hand side of (4.2) has a considerably larger number of multiplications than the logarithmic order on the right-hand side.
Remark 4.1.
The Neumann series can be used to approximate the inverse of matrices for which there exists $\alpha\in\mathbb{R}$ such that $\left\lVert\operatorname{I}-\alpha A\right\rVert_{2}<1$. In that case,
$A^{-1}=\alpha(\operatorname{I}_{n}-(\operatorname{I}_{n}-\alpha A))^{-1}=\alpha\sum_{k=0}^{\infty}(\operatorname{I}_{n}-\alpha A)^{k}.$
The first step is to define a MNN that approximates the function of squaring a matrix $N$ times:
Lemma 4.2.
Suppose that Hypothesis 1 is satisfied, and let $\varepsilon\in(0,\frac{1}{4})$ and $n,N\in\mathbb{N}$. We define $\Phi^{\operatorname{DUP}_{n}}$ as the shallow MNN satisfying:
$R\left(\Phi^{\operatorname{DUP}_{n}}\right)(A)=A|A,$
and $\Phi_{\varepsilon}^{\operatorname{SQR}_{n,\varrho}^{N}}$ as the $\varrho$-MNN satisfying:
$\Phi_{\varepsilon}^{\operatorname{SQR}_{n,\varrho}^{N}}=\overbrace{\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\circ\Phi^{\operatorname{DUP}_{n}}\right)\circ\dots\circ\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\circ\Phi^{\operatorname{DUP}_{n}}\right)}^{N}.$ (4.3)
Then,
$\left\lVert A^{2^{N}}-R\left(\Phi_{\varepsilon}^{\operatorname{SQR}_{n,\varrho}^{N}}\right)(A)\right\rVert_{2}\leq\varepsilon$ (4.4)
for all $A\in\mathbb{R}^{n\times n}$ such that $\left\lVert A\right\rVert_{2}\leq\frac{1}{2}$.
Remark 4.3.
If we define instead the above MNN as
$\Phi_{\varepsilon}^{\operatorname{SQR}_{n,\varrho}^{N}}=\overbrace{\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},\varepsilon+\frac{1}{2^{2^{N-1}}}}\circ\Phi^{\operatorname{DUP}_{n}}\right)\circ\dots\circ\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},\varepsilon+\frac{1}{2^{2}}}\circ\Phi^{\operatorname{DUP}_{n}}\right)}^{N-1}\circ\Phi_{\frac{\varepsilon}{n},\frac{1}{2}}^{\operatorname{STR}_{n,\varrho}}\circ\Phi^{\operatorname{DUP}_{n}},$
it satisfies the same property. This can be seen by analyzing in detail the computations of the proof of Lemma 4.2. However, for clarity and coherence with the result needed for Theorem 4.5 below, we keep the subindices homogeneous and the second one equal to $1$.
Proof.
We can prove (4.4) by induction on $N$. We fix $A\in\mathbb{R}^{n\times n}$ with $\left\lVert A\right\rVert_{2}\leq\frac{1}{2}$.
The case $N=1$ is obvious using (3.16) and (2.1). In fact, we have:
$\left\lVert R\left(\Phi_{\varepsilon}^{\operatorname{SQR}_{n,\varrho}^{1}}\right)(A)-A^{2}\right\rVert_{2}$ $=\left\lVert R\left(\Phi_{\frac{\varepsilon}{4n},1}^{\operatorname{STR}_{n,\varrho}}\right)(A|A)-A^{2}\right\rVert_{2}$ (4.5)
$\leq n\left\lvert R\left(\Phi_{\frac{\varepsilon}{4n},1}^{\operatorname{STR}_{n,\varrho}}\right)(A|A)-A^{2}\right\rvert_{\infty}\leq n\frac{\varepsilon}{4n}=\frac{\varepsilon}{4}<\varepsilon.$
Suppose now that the statement (4.4) holds for $N\in\mathbb{N}$ and let us prove the case $N+1$. By (4.3), we know that:
$\Phi_{\varepsilon}^{\operatorname{SQR}^{N+1}_{n,\varrho}}=\Phi_{\varepsilon}^{\operatorname{SQR}^{1}_{n,\varrho}}\circ\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}.$
By the induction hypothesis, we obtain that:
$\left\lVert R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}\right)(A)\right\rVert_{2}\leq\varepsilon+\left\lVert A\right\rVert_{2}^{2^{N}}\leq\varepsilon+\frac{1}{2^{2^{N}}}\leq\frac{1}{2}.$ (4.6)
Consequently,
$\left\lVert A^{2^{N+1}}-R\left(\Phi_{\varepsilon}^{\operatorname{STR}^{N+1}_{n,\varrho}}\right)(A)\right\rVert_{2}$
$\leq\left\lVert\left(A^{2^{N}}\right)^{2}-\left(R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}\right)(A)\right)^{2}\right\rVert_{2}$
$+\left\lVert\left(R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}\right)(A)\right)^{2}-R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{N+1}_{n,\varrho}}\right)(A)\right\rVert_{2}$
$=\left\lVert A^{2^{N}}\left[A^{2^{N}}-R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}\right)(A)\right]+\left[A^{2^{N}}-R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}\right)(A)\right]R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}\right)(A)\right\rVert_{2}$
$+\left\lVert\left(R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}\right)(A)\right)^{2}-R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{1}_{n,\varrho}}\right)\left(R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}\right)(A)\right)\right\rVert_{2}$
$\leq\left\lVert A^{2^{N}}-R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}\right)(A)\right\rVert_{2}\left(\left\lVert A^{2^{N}}\right\rVert_{2}+\left\lVert R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}\right)(A)\right\rVert_{2}\right)+\frac{\varepsilon}{4}$
$\leq\varepsilon\left(\left\lVert A\right\rVert_{2}^{2^{N}}+\left\lVert R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}\right)(A)\right\rVert_{2}\right)+\frac{\varepsilon}{4}$
$\leq\varepsilon\left(\frac{1}{2^{2^{N}}}+\frac{1}{2}\right)+\frac{\varepsilon}{4}\leq\varepsilon,$
where we used the inductive hypothesis (4.4) and also used (4.5) since (4.6) holds. ∎
Remark 4.4.
This Matrix Neural Network can be used to compute any power of a matrix. Using the binary representation of the exponent, any power of a matrix can be expressed as the product of powers of 2 of that matrix. In other words, if we want to calculate $A^{M}$ with $A\in\mathbb{R}^{n\times n}$ and $M\in\mathbb{R}$, then, using the fact that there exist $N_{1},\dots,N_{k}\in\mathbb{N}\cup\left\{0\right\}$ all different such that $M=2^{N_{1}}+\dots+2^{N_{k}}$, $A^{M}=A^{2^{N_{1}}}\cdots A^{2^{N_{k}}}$. Since the product of matrices and the raising to a powers of 2 can be done with Matrix Neural Networks, the operation of raising a matrix to an arbitrary exponent can be expressed as such too.
Theorem 4.5.
Suppose that Hypothesis 1 is satisfied and let $\varepsilon\in(0,\frac{1}{8})$ and $n,N\in\mathbb{N}$. Then, if $N\geq 2$ there exists a $\varrho$-MNN $\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{N}}$ satisfying:
$\left\lVert\sum_{k=0}^{2^{N}-1}A^{k}-R\left(\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{N}}\right)(A)\right\rVert_{2}\leq\varepsilon$ (4.7)
for all $A\in\mathbb{R}^{n\times n}$ such that $\left\lVert A\right\rVert_{2}\leq 1$, satisyfing:
$\begin{split}M\left(\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{N}}\right)&\leq 14n^{\log_{2}7}(N-1)\left(M\left(\Pi_{2^{-2^{N}}\frac{\varepsilon}{8n^{3}},2n}^{\varrho}\right)+12\right)\\ &+n^{2}\left(L\left(\Pi_{2^{-2^{N}}\frac{\varepsilon}{8n^{3}},2n}^{\varrho}\right)-14N+19+2\log_{2}n\right)+nN,\end{split}$ (4.8)
and satisfying:
$L\left(\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{N}}\right)\leq N\left[2\log_{2}n+5+L\left(\Pi_{2^{-2^{N}}\frac{\varepsilon}{8n^{3}},2n}^{\varrho}\right)\right].$ (4.9)
Also, if $N=1$ there exists a $\varrho$-MNN $\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{N}}$ satisfying (4.7) for $\varepsilon=0$ and all $A\in\mathbb{R}^{n\times n}$, and satisfying:
$M\left(\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{1}}\right)$ $=n^{2}+n,$ (4.10)
$L\left(\Phi_{\varepsilon,K}^{\operatorname{NEU}_{n,\varrho}^{1}}\right)$ $=1.$ (4.11)
Part of the improvements with respect to the paper [KPRS22] can be found in the Master Thesis [Rom24].
Proof.
We split the proof into 6 steps:
  • Step 1: The simple case $N=1$.
  • Step 2: Informal definition of the MNN.
  • Step 3: Formal definition of the MNN.
  • Step 4: Proof of (4.7) for $N\geq 2$.
  • Step 5: Proof of (4.8) for $N\geq 2$.
  • Step 6: Proof of (4.9) for $N\geq 2$.
Step 1: The simple case $N=1$.
We define:
$\Phi^{\operatorname{NEU}^{1}_{n,\varrho}}_{\varepsilon}=((\operatorname{id},\operatorname{I}_{n},\operatorname{id})).$
Then, it is immediate that $R\left(\Phi^{\operatorname{NEU}^{1}_{n,\varrho}}_{\varepsilon}\right)(A)=A+\operatorname{I}_{n}$ with $M\left(\Phi^{\operatorname{NEU}^{1}_{n,\varrho}}_{\varepsilon}\right)=n^{2}+n$ and $L\left(\Phi^{\operatorname{NEU}^{1}_{n,\varrho}}_{\varepsilon}\right)=1$.
Step 2: Informal definition of the MNN.
We now suppose that $N\geq 2$.
The general idea is to use the product formula introduced in (4.2) for the Neumann series:
$\sum_{k=0}^{2^{N+1}-1}A^{k}=\prod_{k=0}^{N}\left(A^{2^{k}}+\operatorname{I}_{n}\right).$
However, by our construction of $\Phi^{\operatorname{STR}}$, we require to bound the norm of the approximation of $\operatorname{I}_{n}+A^{2^{k}}$ and $A^{2^{k}}$ for $k\in\left\{0,\dots,N\right\}$ for an accurate product. For that purpose, we rewrite the previous product as follows:
$\begin{split}\prod_{k=0}^{N}\left(A^{2^{k}}+\operatorname{I}_{n}\right)&=2^{\sum_{k=0}^{N}2^{k}}\prod_{k=0}^{N}2^{-2^{k}}\left(A^{2^{k}}+\operatorname{I}_{n}\right)\\ &=2^{2^{N+1}-1}\prod_{k=0}^{N}\left[\left(\frac{A}{2}\right)^{2^{k}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{k}}\right].\end{split}$ (4.12)
We do this to have upper bounds of the approximations with the cost of the power of two as a factor.
The target MNN will then have two branches: one branch responsible for approximating $\left(\frac{A}{2}\right)^{2^{k}}$ for each $k\in\left\{0,\dots,N\right\}$ like in Lemma 4.2, while the other approximates the target product using recursively the different exponents of $\frac{A}{2}$. Our plan is then to follow the sketch:
$A$ $\mapsto\begin{pmatrix}\frac{A}{2}&|&\frac{A}{2}\\ \frac{A}{2}&|&0\end{pmatrix}\mapsto\begin{pmatrix}\left(\frac{A}{2}\right)^{2}\\ \frac{A}{2}+\frac{\operatorname{I}_{n}}{2}\end{pmatrix}$ (4.13)
$\mapsto\begin{pmatrix}\left(\frac{A}{2}\right)^{2}&|&\left(\frac{A}{2}\right)^{2}\\ \left(\frac{A}{2}\right)^{2}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2}&|&\frac{A}{2}+\frac{\operatorname{I}_{n}}{2}\end{pmatrix}\mapsto\begin{pmatrix}\left(\frac{A}{2}\right)^{4}\\ \left(\frac{A}{2}+\frac{\operatorname{I}_{n}}{2}\right)\left(\left(\frac{A}{2}\right)^{2}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2}\right)\end{pmatrix}\mapsto\cdots$
$\mapsto\begin{pmatrix}\left(\frac{A}{2}\right)^{2^{N-1}}&|&\left(\frac{A}{2}\right)^{2^{N-1}}\\ \left(\frac{A}{2}\right)^{2^{N-1}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{N-1}}&|&\prod_{k=0}^{N-2}\left(\left(\frac{A}{2}\right)^{2^{k}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{k}}\right)\end{pmatrix}$
$\mapsto\begin{pmatrix}\left(\frac{A}{2}\right)^{2^{N}}\\ \prod_{k=0}^{N-1}\left(\left(\frac{A}{2}\right)^{2^{k}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{k}}\right)\end{pmatrix}$
$\mapsto\left(\frac{A}{2}\right)^{2^{N}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{N}}\bigg{|}\prod_{k=0}^{N-1}\left(\left(\frac{A}{2}\right)^{2^{k}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{k}}\right)$
$\mapsto 2^{2^{N+1}-1}\prod_{k=0}^{N}\left(\left(\frac{A}{2}\right)^{2^{k}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{k}}\right).$
This way we compute $\left(\frac{A}{2}\right)^{2^{k}}$ only once for every $k\in\mathbb{N}$.
Step 3: Formal definition of the MNN.
We define some auxiliary MNN satisfying:
$R\left(\Phi^{\operatorname{DUP}}_{n}\right)(A)$ $=\frac{1}{2}\begin{pmatrix}A&|&A\\ A&|&0\end{pmatrix},$
$R\left(\Phi^{\operatorname{FILL}}_{n,L}\right)(A|B)$ $=A+\frac{\operatorname{I}_{n}}{2},$
$R\left(\Phi^{\operatorname{FLIP}}_{n,k}\right)\begin{pmatrix}A\\ B\end{pmatrix}$ $=(A+2^{-2^{k}}\operatorname{I}_{n})|B,$
$R\left(\Phi^{\operatorname{MIX}}_{n,k}\right)\begin{pmatrix}A\\ B\end{pmatrix}$ $=\begin{pmatrix}A&|&A\\ A+2^{-2^{k}}\operatorname{I}_{n}&|&B\end{pmatrix},$
for all $A,B\in\mathbb{R}^{n\times n}$, $k,L\in\mathbb{N}$ with
$M\left(\Phi^{\operatorname{DUP}}_{n}\right)$ $=3n^{2},$ $L\left(\Phi^{\operatorname{DUP}}_{n}\right)$ $=1,$ (4.14)
$M\left(\Phi^{\operatorname{FILL}}_{n,L}\right)$ $=n^{2}L+n,$ $L\left(\Phi^{\operatorname{FILL}}_{n,L}\right)$ $=L,$
$M\left(\Phi^{\operatorname{FLIP}}_{n,k}\right)$ $=2n^{2}+n,$ $L\left(\Phi^{\operatorname{FLIP}}_{n,k}\right)$ $=1,$
$M\left(\Phi^{\operatorname{MIX}}_{n,k}\right)$ $=4n^{2}+n,$ $L\left(\Phi^{\operatorname{MIX}}_{n,k}\right)$ $=1.$
The MNN $\Phi_{n,L}^{\operatorname{FILL}}$ is artificially elongated in order to parallelize it with a MNN with more than 1 layer. We can then define recursively that:
$\begin{cases}\Phi^{\operatorname{AUX}^{1}_{n,\varrho}}_{\varepsilon}=\mathcal{P}\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1},\Phi^{\operatorname{FILL}}_{n,L\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\right)}\right)\circ\Phi^{\operatorname{DUP}}_{n},\\ \Phi^{\operatorname{AUX}^{i}_{n,\varrho}}_{\varepsilon}=\mathcal{P}\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1},\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\right)\circ\Phi^{\operatorname{MIX}}_{n,i-1}\circ\Phi^{\operatorname{AUX}^{i-1}_{n,\varrho}}_{\varepsilon}&\forall i\geq 2.\end{cases}$ (4.15)
If we define:
$\left((\mathcal{L}^{1},C^{1},\alpha^{1}),\dots,(\mathcal{L}^{L},C^{L},\alpha^{L})\right)\coloneq\Phi^{\operatorname{STR}_{n,\varrho}}_{2^{1-2^{N}}\frac{\varepsilon}{4n},1}\circ\Phi^{\operatorname{FLIP}}_{n,N-1}\circ\Phi^{\operatorname{AUX}^{N-1}_{n,\varrho}}_{2^{1-2^{N}}\varepsilon},$
where we have replaced $\varepsilon$ by $2^{1-2^{N}}\varepsilon$, then the target $\varrho$-MNN is:
$\Phi_{\varepsilon}^{\operatorname{NEU}^{N}_{n,\varrho}}\coloneq\left((\mathcal{L}^{1},C^{1},\alpha^{1}),\dots,\left(2^{2^{N}-1}\mathcal{L}^{L},2^{2^{N}-1}C^{L},\alpha^{L}\right)\right).$
For legibility, we introduce some notation useful for the next steps:
$\operatorname{SQR}_{\varepsilon}^{i}(A)$ $=\left[R\left(\Phi_{\varepsilon}^{\operatorname{AUX}^{i}_{n,\varrho}}\right)(A)\right]_{1:n,1:n},$
$\operatorname{PROD}_{\varepsilon}^{i}(A)$ $=\left[R\left(\Phi_{\varepsilon}^{\operatorname{AUX}^{i}_{n,\varrho}}\right)(A)\right]_{(n+1):2n,1:n},$
$P^{i}(A)$ $=\prod_{k=0}^{i}\left(\left(\frac{A}{2}\right)^{2^{k}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{k}}\right)$
for all $i\in\mathbb{N}$ and $A\in\mathbb{R}^{n\times n}$. It is not difficult to see using (4.3) and (4.15) that:
$\operatorname{SQR}^{i}_{\varepsilon}(A)=R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{i}_{n,\varrho}}\right)\left(\frac{A}{2}\right),$ (4.16)
where $\Phi_{\varepsilon}^{\operatorname{SQR}^{i}_{n,\varrho}}$ is the $\varrho$-MNN from Lemma 4.2. Thus, we know by (4.4) that:
$\left\lVert\operatorname{SQR}^{i}_{\varepsilon}(A)\right\rVert_{2}\leq\varepsilon+\left\lVert\frac{A}{2}\right\rVert_{2}^{2^{i}}\leq\varepsilon+2^{-2^{i}}$ (4.17)
for all $i\in\mathbb{N}$ and $A\in\mathbb{R}^{n\times n}$ such that $\left\lVert A\right\rVert_{2}\leq 1$. It is also easy to check that:
$\left\lVert P^{i}(A)\right\rVert_{2}\leq\prod_{k=0}^{i}\frac{2}{2^{2^{k}}}=2^{i+2-2^{i+1}}\leq\frac{1}{2}$ (4.18)
for all $i\in\mathbb{N}$ and $A\in\mathbb{R}^{n\times n}$ with $\left\lVert A\right\rVert_{2}\leq 1$.
[…]
$\operatorname{SQR}^{i-1}_{\varepsilon}(A)+\frac{\operatorname{I}_{n}}{2^{2^{i-1}}}\Big{|}\operatorname{PROD}^{i-1}_{\varepsilon}(A)$
[…]
Figure 1: Matrix Neural Network scheme of $\Phi^{\operatorname{AUX}^{i}_{n,\varrho}}_{\varepsilon}$ for all $i\geq 2$.
Figure 1 visualizes the process that follows an input matrix $A$ in the MNN $\Phi^{\operatorname{AUX}^{i}_{n,\varrho}}_{\varepsilon}$. Basically, the operations that appear in the informal sketch (4.13) are replaced by the approximations performed by the MNNs introduced previously.
Step 4: Proof of (4.7) for $N\geq 2$.
Equations (4.4) and (4.16) give us
$\left\lVert\left(\frac{A}{2}\right)^{2^{i}}-\operatorname{SQR}^{i}_{\varepsilon}(A)\right\rVert_{2}\leq\varepsilon$
for all $i\in\mathbb{N}$ and $A\in\mathbb{R}^{n\times n}$ such that $\left\lVert A\right\rVert_{2}\leq 1$.
Let us see by induction over $i\geq 2$ that:
$\left\lVert P^{i-1}(A)-\left[R\left(\Phi^{\operatorname{AUX}^{i}_{n,\varrho}}_{\varepsilon}\right)(A)\right]_{n+1:2n,1:n}\right\rVert_{2}\leq\varepsilon$ (4.19)
for all $A\in\mathbb{R}^{n\times n}$ such that $\left\lVert A\right\rVert_{2}\leq 1$.
We start with the base case $i=2$. Let $A\in\mathbb{R}^{n\times n}$ be such that $\left\lVert A\right\rVert_{2}\leq 1$. We get then
$\left\lVert P^{1}(A)-\operatorname{PROD}_{\varepsilon}^{2}(A)\right\rVert_{2}$ $=\left\lVert\frac{A+\operatorname{I}_{n}}{2}\left(\left(\frac{A}{2}\right)^{2}+\frac{\operatorname{I}_{n}}{4}\right)\right.$
$\left.-R\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\right)\left(\operatorname{SQR}^{1}_{\varepsilon}(A)+\frac{\operatorname{I}_{n}}{4}\bigg{|}\frac{A+\operatorname{I}_{n}}{2}\right)\right\rVert_{2}$
$\leq\left\lVert\left(\left(\frac{A}{2}\right)^{2}-\operatorname{SQR}_{\varepsilon}^{1}(A)\right)\left(\frac{A+\operatorname{I}_{n}}{2}\right)\right\rVert_{2}$
$+\left\lVert\left(\operatorname{SQR}_{\varepsilon}^{1}(A)+2^{-2}\operatorname{I}_{n}\right)\left(\frac{A+\operatorname{I}_{n}}{2}\right)\right.$
$\left.-R\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\right)\left(\operatorname{SQR}^{1}_{\varepsilon}(A)+2^{-2}\operatorname{I}_{n}\bigg{|}\frac{A+\operatorname{I}_{n}}{2}\right)\right\rVert_{2}$
$=\left\lVert\left(\left(\frac{A}{2}\right)^{2}-R\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\right)\left(\frac{A}{2}\bigg{|}\frac{A}{2}\right)\right)\left(\frac{A+\operatorname{I}_{n}}{2}\right)\right\rVert_{2}$
$+\left\lVert\left(\operatorname{SQR}_{\varepsilon}^{1}(A)+2^{-2}\operatorname{I}_{n}\right)\left(\frac{A+\operatorname{I}_{n}}{2}\right)\right.$
$\left.-R\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\right)\left(\operatorname{SQR}^{1}_{\varepsilon}(A)+2^{-2}\operatorname{I}_{n}\bigg{|}\frac{A+\operatorname{I}_{n}}{2}\right)\right\rVert_{2}$
$\leq\frac{\varepsilon}{4}+\frac{\varepsilon}{4}\leq\varepsilon,$
where we have used that
$\left\lVert\frac{A+\operatorname{I}_{n}}{2}\right\rVert_{2}\leq 1,$
that
$\left\lVert\operatorname{SQR}^{1}_{\varepsilon}(A)+2^{-2}\operatorname{I}_{n}\right\rVert_{2}\leq 1,$
as well as (4.4) and (2.1).
Let us suppose that the result (4.19) is true for $i\geq 2$ and prove it for $i+1$. Using (4.18) and the inductive hypothesis (4.19), we get that
$\left\lVert\operatorname{PROD}^{i}_{\varepsilon}(A)\right\rVert_{2}\leq\varepsilon+\left\lVert P^{i}(A)\right\rVert_{2}\leq\varepsilon+\frac{1}{2}<1.$
Consequently, using (2.1) and (4.17) we get:
$\left\lVert\left(\operatorname{SQR}_{\varepsilon}^{i}(A)+2^{-2^{i}}\operatorname{I}_{n}\right)\operatorname{PROD}_{\varepsilon}^{i}(A)\right.\\ \left.-R\left(\Phi_{\frac{\varepsilon}{4n},1}^{\operatorname{STR}_{n,\varrho}}\right)\left(\operatorname{SQR}_{\varepsilon}^{i}(A)+2^{-2^{i}}\operatorname{I}_{n}\middle|\operatorname{PROD}_{\varepsilon}^{i}(A)\right)\right\rVert_{2}\leq\frac{\varepsilon}{4}.$
We conclude then
$\left\lVert P^{i}(A)-\operatorname{PROD}_{\varepsilon}^{i+1}(A)\right\rVert_{2}=\left\lVert\left(2^{-2^{i}}\operatorname{I}_{n}+\left(\frac{A}{2}\right)^{2^{i}}\right)P^{i-1}(A)\right.$
$\left.-R\left(\Phi_{\frac{\varepsilon}{4n},1}^{\operatorname{STR}_{n,\varrho}}\right)\left(\operatorname{SQR}_{\varepsilon}^{i}(A)+2^{-2^{i}}\operatorname{I}_{n}\middle|\operatorname{PROD}_{\varepsilon}^{i}(A)\right)\right\rVert_{2}$
$\leq\left\lVert\left(2^{-2^{i}}\operatorname{I}_{n}+\left(\frac{A}{2}\right)^{2^{i}}\right)P^{i-1}(A)-\left(\operatorname{SQR}_{\varepsilon}^{i}(A)+2^{-2^{i}}\operatorname{I}_{n}\right)\operatorname{PROD}_{\varepsilon}^{i}(A)\right\rVert_{2}+\frac{\varepsilon}{4}$
$\leq\left\lVert\left(\left(\frac{A}{2}\right)^{2^{i}}-\operatorname{SQR}_{\varepsilon}^{i}(A)\right)P^{i-1}(A)\right\rVert_{2}$
$+\left\lVert\left(2^{-2^{i}}\operatorname{I}_{n}+\operatorname{SQR}^{i}_{\varepsilon}(A)\right)\left(P^{i-1}(A)-\operatorname{PROD}_{\varepsilon}^{i}(A)\right)\right\rVert_{2}+\frac{\varepsilon}{4}$
$\leq\varepsilon\left(\left\lVert P^{i-1}(A)\right\rVert_{2}+\left\lVert 2^{-2^{i}}\operatorname{I}_{n}+\operatorname{SQR}_{\varepsilon}^{i}(A)\right\rVert_{2}+\frac{1}{4}\right)$
$\leq\varepsilon\left(2^{i-2^{i}}+2^{-2^{i}}+\varepsilon+2^{-2^{i}}+\frac{1}{4}\right)\leq\varepsilon,$
which implies (4.19). This also implies (4.7) because for all $A\in\mathbb{R}^{n\times n}$:
$R\left(\Phi^{\operatorname{NEU}^{N}_{n,\varrho}}_{\varepsilon}\right)(A)$
$=2^{2^{N}-1}R\left(\Phi_{2^{1-2^{N}}\frac{\varepsilon}{4n},1}^{\operatorname{STR}_{n,\varrho}}\right)\left(\operatorname{SQR}^{N-1}_{2^{1-2^{N}}\varepsilon}(A)+2^{-2^{N-1}}\operatorname{I}_{n}\middle|\operatorname{PROD}_{2^{1-2^{N}}\varepsilon}^{N-1}(A)\right)$
$=\operatorname{PROD}_{\varepsilon}^{N}(A).$
where we used the definitions of $\Phi_{\varepsilon}^{\operatorname{NEU}^{N}_{n,\varrho}}$, $\operatorname{PROD}_{2^{-2^{N}}\varepsilon}^{N}$, and $\Phi^{\operatorname{FLIP}}_{n,N-1}$.
Step 5: Proof of (4.8) for $N\geq 2$.
We can rapidly deduce with (4.14) that
$M\left(\Phi_{\varepsilon}^{\operatorname{NEU}^{N}_{n,\varrho}}\right)$ $=M\left(\Phi_{2^{1-2^{N}}\frac{\varepsilon}{4n},1}^{\operatorname{\operatorname{STR}_{n,\varrho}}}\right)+M\left(\Phi^{\operatorname{FLIP}}_{n,N-1}\right)$
$+\sum_{k=2}^{N-1}\left[M\left(\mathcal{P}\left(\Phi_{2^{1-2^{N}}\frac{\varepsilon}{4n},1}^{\operatorname{\operatorname{STR}_{n,\varrho}}},\Phi_{2^{1-2^{N}}\frac{\varepsilon}{4n},1}^{\operatorname{\operatorname{STR}_{n,\varrho}}}\right)\right)+M\left(\Phi_{n,k-1}^{\operatorname{MIX}}\right)\right]$
$+M\left(\mathcal{P}\left(\Phi_{2^{1-2^{N}}\frac{\varepsilon}{4n},1}^{\operatorname{\operatorname{STR}_{n,\varrho}}},\Phi_{n,L\left(\Phi_{2^{1-2^{N}}\frac{\varepsilon}{4n},1}^{\operatorname{\operatorname{STR}_{n,\varrho}}}\right)}^{\operatorname{FILL}}\right)\right)+M\left(\Phi^{\operatorname{DUP}}_{n}\right)$
$=2(N-1)M\left(\Phi_{2^{-2^{N}}\frac{\varepsilon}{2n},1}^{\operatorname{\operatorname{STR}_{n,\varrho}}}\right)+2n^{2}+n+(N-2)(4n^{2}+n)$
$+n^{2}L\left(\Phi_{2^{-2^{N}}\frac{\varepsilon}{2n},1}^{\operatorname{\operatorname{STR}_{n,\varrho}}}\right)+n+3n^{2}.$
Thus, using (3.17) and (3.18), we obtain that:
$M\left(\Phi_{\varepsilon}^{\operatorname{NEU}^{N}_{n,\varrho}}\right)$ $\leq 2(N-1)\left[7n^{\log_{2}7}\left(M\left(\Pi_{2^{1-2^{N}}\frac{\varepsilon}{16n^{3}},2n}^{\varrho}\right)+12\right)-9n^{2}\right]$
$+n^{2}\left[4N-3+L\left(\Pi_{2^{1-2^{N}}\frac{\varepsilon}{16n^{3}},2n}^{\varrho}\right)+2(\log_{2}n+2)\right]+nN$
$=14n^{\log_{2}7}(N-1)\left(M\left(\Pi_{2^{-2^{N}}\frac{\varepsilon}{8n^{3}},2n}^{\varrho}\right)+12\right)$
$+n^{2}\left(L\left(\Pi_{2^{-2^{N}}\frac{\varepsilon}{8n^{3}},2n}^{\varrho}\right)-14N+19+2\log_{2}n\right)+nN.$
Step 6: Proof of (4.9) for $N\geq 2$.
For the number of layers (4.9), we can similarly to Step 5 get that
$L\left(\Phi_{\varepsilon}^{\operatorname{NEU}^{N}_{n,\varrho}}\right)$ $=N\left(L\left(\Phi_{2^{1-2^{N}}\frac{\varepsilon}{4n},1}^{\operatorname{\operatorname{STR}_{n,\varrho}}}\right)+1\right)$
$\leq N\left[2\log_{2}n+5+L\left(\Pi_{2^{-2^{N}}\frac{\varepsilon}{8n^{3}},2n}^{\varrho}\right)\right],$
using again (4.14) and (3.18). ∎
We can finally get a MNN approximating the inverse of some matrices.
Theorem 4.6.
Suppose that Hypothesis 1 is satisfied. Let $\varepsilon>0$, $\alpha\in\mathbb{R}$, $\delta\in(0,1)$ and $n\in\mathbb{N}$. We call
$N(\varepsilon,\delta)=\left\lceil\log_{2}\left(\max\left\{\frac{\log_{2}(\varepsilon(1-\delta))}{\log_{2}\delta},2\right\}\right)\right\rceil.$ (4.20)
Then, there exists a $\varrho$-MNN $\Phi^{\operatorname{INV}_{n,\varrho}^{\alpha}}_{\varepsilon,\delta}$ satisfying:
$\left\lVert A^{-1}-R\left(\Phi^{\operatorname{INV}_{n,\varrho}^{\alpha}}_{\varepsilon,\delta}\right)(A)\right\rVert_{2}\leq\varepsilon$ (4.21)
for all $A\in\mathbb{R}^{n\times n}$ such that
$\left\lVert\operatorname{I}_{n}-\alpha A\right\rVert_{2}\leq\delta,$
and satisfying:
$M\left(\Phi^{\operatorname{INV}_{n,\varrho}^{\alpha}}_{\varepsilon,\delta}\right)$ $\leq 14n^{\log_{2}7}\left(N\left(\frac{\varepsilon}{2\alpha},\delta\right)-1\right)\left(M\left(\Pi^{\varrho}_{2^{-2^{N\left(\frac{\varepsilon}{2\alpha},\delta\right)}}\frac{\min\left\{\frac{\varepsilon}{\alpha},\frac{1}{4}\right\}}{16n^{3}},2n}\right)+12\right)$
$+n^{2}\left(L\left(\Pi_{2^{-2^{N\left(\frac{\varepsilon}{2\alpha},\delta\right)}}\frac{\min\left\{\frac{\varepsilon}{\alpha},\frac{1}{4}\right\}}{16n^{3}},2n}^{\varrho}\right)-14N\left(\frac{\varepsilon}{2\alpha},\delta\right)+20+2\log_{2}n\right)$ (4.22)
$+n\left(N\left(\frac{\varepsilon}{2\alpha},\delta\right)+1\right),$
$L\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{\alpha}_{n,\varrho}}\right)$ $\leq N\left(\frac{\varepsilon}{2\alpha},\delta\right)\left[2\log_{2}n+5+L\left(\Pi_{2^{-2^{N\left(\frac{\varepsilon}{2\alpha},\delta\right)}}\frac{\min\left\{\frac{\varepsilon}{\alpha},\frac{1}{4}\right\}}{16n^{3}},2n}^{\varrho}\right)\right],$ (4.23)
if $N\left(\frac{\varepsilon}{2\alpha},\delta\right)\geq 2$, and satisfying:
$M\left(\Phi^{\operatorname{INV}_{n,\varrho}^{\alpha}}_{\varepsilon,\delta}\right)$ $=2(n^{2}+n),$ (4.24)
$L\left(\Phi^{\operatorname{INV}_{n,\varrho}^{\alpha}}_{\varepsilon,\delta}\right)$ $=2,$ (4.25)
if $N\left(\frac{\varepsilon}{2\alpha},\delta\right)=1$.
Note that $\varepsilon\mapsto N(\varepsilon,\delta)$ decreases with $\varepsilon$ when $\delta\in(0,1)$.
Proof.
We first focus on proving (4.21).
We know that the identity
$A^{-1}=\alpha\sum_{k=0}^{\infty}(\operatorname{I}_{n}-\alpha A)^{k}$
holds for all $A\in\mathbb{R}^{n\times n}$ such that $\left\lVert\operatorname{I}_{n}-\alpha A\right\rVert_{2}<1$. By (4.1), we have that
$\begin{split}\left\lVert A^{-1}-\alpha\sum_{k=0}^{2^{N}-1}(\operatorname{I}_{n}-\alpha A)^{k}\right\rVert_{2}&=\alpha\left\lVert(\operatorname{I}_{n}-(\operatorname{I}_{n}-\alpha A))^{-1}-\sum_{k=0}^{2^{N}-1}(\operatorname{I}_{n}-\alpha A)^{k}\right\rVert\\ &\leq\alpha\frac{\left\lVert\operatorname{I}_{n}-\alpha A\right\rVert_{2}^{2^{N}}}{1-\left\lVert\operatorname{I}_{n}-\alpha A\right\rVert_{2}}\leq\alpha\frac{\delta^{2^{N}}}{1-\delta}\end{split}$ (4.26)
when
$\left\lVert\operatorname{I}_{n}-\alpha A\right\rVert_{2}\leq\delta\in(0,1).$
Our plan is then to use the MNN in Theorem 4.5 to approximate matrix inversion. It is easy to check that for $N(\varepsilon,\delta)$ given by (4.20) we get the inequality:
$\frac{\delta^{2^{N(\varepsilon,\delta)}}}{1-\delta}\leq\varepsilon.$ (4.27)
Consider the shallow MNN $\Phi^{\operatorname{IN}}_{n,\alpha}$ that satisfies:
$R\left(\Phi^{\operatorname{IN}}_{n,\alpha}\right)(A)=\operatorname{I}_{n}-\alpha A$
for all $A\in\mathbb{R}^{n\times n}$ and
$M\left(\Phi^{\operatorname{IN}}_{n}\right)=n+n^{2}.$ (4.28)
If we define
$((\mathcal{L}^{1},B^{1},\alpha^{1}),\dots,(\mathcal{L}^{L},B^{L},\alpha^{L}))\coloneq\Phi_{\min\left\{\frac{\varepsilon}{2\alpha},\frac{1}{8}\right\}}^{\operatorname{NEU}^{N\left(\frac{\varepsilon}{2\alpha},\delta\right)}_{n,\varrho}},$
then our target MNN is
$\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{\alpha}_{n,\varrho}}\coloneq((\mathcal{L}^{1},B^{1},\alpha^{1}),\dots,(\alpha\mathcal{L}^{L},\alpha B^{L},\alpha^{L}))\circ\Phi_{n,\alpha}^{\operatorname{IN}}.$
Indeed, we get for all $A\in\mathbb{R}^{n\times n}$ such that $\left\lVert\operatorname{I}_{n}-\alpha A\right\rVert_{2}\leq\delta$:
$\left\lVert A^{-1}-R\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{\alpha}_{n,\varrho}}\right)(A)\right\rVert_{2}$
$\leq\left\lVert A^{-1}-\alpha\sum^{2^{N\left(\frac{\varepsilon}{2\alpha},\delta\right)}-1}_{k=0}(\operatorname{I}_{n}-\alpha A)^{k}\right\rVert_{2}$
$+\left\lVert\alpha\sum^{2^{N\left(\frac{\varepsilon}{2\alpha},\delta\right)}-1}_{k=0}(\operatorname{I}_{n}-\alpha A)^{k}-\alpha R\left(\Phi_{\min\left\{\frac{\varepsilon}{2\alpha},\frac{1}{8}\right\}}^{\operatorname{NEU}^{N\left(\frac{\varepsilon}{2\alpha},\delta\right)}_{n,\varrho}}\right)(\operatorname{I}_{n}-\alpha A)\right\rVert_{2}$
$\leq\alpha\frac{\delta^{2^{N\left(\frac{\varepsilon}{2\alpha},\delta\right)}}}{1-\delta}+\alpha\min\left\{\frac{\varepsilon}{2\alpha},\frac{1}{8}\right\}\leq\varepsilon.$
In the second estimate, we have used (4.26) and (4.7), and in the third one (4.27).
We follow by proving the upper bound in the number of weights. For that, we consider that:
$M\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{\alpha}_{n,\varrho}}\right)=M\left(\Phi_{\min\left\{\frac{\varepsilon}{2\alpha},\frac{1}{8}\right\}}^{\operatorname{NEU}^{N\left(\frac{\varepsilon}{2\alpha},\delta\right)}_{n,\varrho}}\right)+M\left(\Phi^{\operatorname{IN}}_{n,\alpha}\right).$ (4.29)
In fact, if $N\left(\frac{\varepsilon}{2\alpha},\delta\right)\geq 2$ we obtain (4.22) from (4.29), (4.8) and (4.28). Similarly, if $N\left(\frac{\varepsilon}{2\alpha},\delta\right)=1$, we obtain (4.24) from (4.29), (4.10) and (4.28).
We finish by proving the upper bound in the number of layers, that is, by proving (4.23) and (4.25). For that, we only need to consider:
$L\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{\alpha}_{n,\varrho}}\right)=L\left(\Phi_{\min\left\{\frac{\varepsilon}{2\alpha},\frac{1}{8}\right\}}^{\operatorname{NEU}^{N\left(\frac{\varepsilon}{2\alpha},\delta\right)}_{n,\varrho}}\right)+1,$
and the estimates (4.9) and (4.11) respectively. ∎
5 Applications for resolution of elliptic
parametric PDEs
As explained in the paper [KPRS22], this way of inverting matrices allows to numerically solve some parametric elliptic PDEs.
Consider the weak form of an elliptic PDE:
$b_{y}(u_{y},v)=f_{y}(v)\quad\text{ for all }v\in H,y\in\mathcal{Y},$ (5.1)
where
  1. 1.
    $H$ is a infinite dimensional separable Hilbert space with $\left\lVert\;\cdot\;\right\rVert_{H}$ its associated norm,
  2. 2.
    $\mathcal{Y}$ is the parameter set,
  3. 3.
    $b_{y}$ is a parameter dependent bilinear form,
  4. 4.
    $f_{y}$ is a parameter dependent linear map in the dual of $H$,
  5. 5.
    $u_{y}\in H$ is the solution of (5.1).
A way to garantee the existence and uniqueness of $u_{y}$ is assuming some properties on the bilinear forms $b_{y}$:
  1. 1.
    $b_{y}$ are symmetric, that is, for all $y\in\mathcal{Y}$ and $u,v\in H$, $b_{y}(u,v)=b_{y}(v,u)$.
  2. 2.
    $b_{y}$ are uniformly continuous, that is, there exists a constant $C_{\operatorname{cont}}>0$ such that
    $\left\lvert b_{y}(u,v)\right\rvert\leq C_{\operatorname{cont}}\left\lVert u\right\rVert_{H}\left\lVert v\right\rVert_{H},\quad\forall y\in\mathcal{Y},\ \ \forall u,v\in H.$
  3. 3.
    $b_{y}$ are uniformly coercive, that is, there exists a constant $C_{\operatorname{coer}}>0$ such that
    $b_{y}(u,u)\geq C_{\operatorname{coer}}\left\lVert u\right\rVert_{H}^{2},\quad\forall y\in\mathcal{Y},\ \ \forall u\in H.$
In fact, the Lax-Milgram Theorem ensures it, as proven in Theorem 1 of Section 6.2 of [Eva10].
In this situation, we can use the Galerkin method to discretize the Hilbert space truncating a Hilbert base to transform the PDE into a linear system of equations. The associated matrix of said system is defined as
$B_{y}=\left(b_{y}(\varphi_{i},\varphi_{j})\right)_{\begin{subarray}{c}i=1,\dots,n\\ j=1,\dots,n\end{subarray}}\in\mathbb{R}^{n\times n},$
where $\left\{\varphi_{i}\right\}_{i\in\mathbb{N}}$ is an orthonormal basis of $H$ and $n\in\mathbb{N}$ is the number of elements of said basis after the truncation. We are going to see, with the assumption on $b_{y}$, that the matrices $B_{y}$ satisfy the requirement that there exist $\alpha\in\mathbb{R}$ and $\delta\in(0,1)$ independent of $y\in\mathcal{Y}$ such that $\left\lVert\operatorname{I}_{n}-\alpha B_{y}\right\rVert_{2}<\delta$. This property makes the MNN of Theorem 4.6 particularly useful: one MNN gives the Galerkin solution for all $y\in\mathcal{Y}$ as long as we are able to obtain the matrices $B_{y}$.
Proposition 5.1.
Let $H$ be a separable Hilbert space with $\left\{\varphi_{i}\right\}_{i\in\mathbb{N}}$ an orthonormal basis, $\mathcal{Y}$ an arbitrary set, and $\left\{b_{y}\right\}_{y\in\mathcal{Y}}$ be a collection of bilinear forms in $H$. Suppose that $\left\{b_{y}\right\}_{y\in\mathcal{Y}}$ are symmetric, uniformly continuous with constant $C_{\operatorname{cont}}$ and uniformly coercive with constant $C_{\operatorname{coer}}$. Then, for every $\alpha\in\left(0,\frac{2}{C_{\operatorname{cont}}}\right)$ and $n\in\mathbb{N}$ all matrices
$B_{y}=\left(b_{y}(\varphi_{i},\varphi_{j})\right)_{\begin{subarray}{c}i=1,\dots,n\\ j=1,\dots,n\end{subarray}}\in\mathbb{R}^{n\times n}$
satisfy:
$\left\lVert\operatorname{I}_{n}-B_{y}\right\rVert_{2}\leq\max\left\{\left\lvert 1-\alpha C_{\operatorname{cont}}\right\rvert,\left\lvert 1-\alpha C_{\operatorname{coer}}\right\rvert\right\}<1.$
Remark 5.2.
This proposition is a generalization of Proposition B.1 of [KPRS22], where they consider $\alpha\in\left(0,\frac{1}{C_{\operatorname{cont}}}\right)$ instead of $\alpha\in\left(0,\frac{2}{C_{\operatorname{cont}}}\right)$.
Proof.
Since $B_{y}$ is symmetric for $y\in\mathcal{Y}$ fixed, it is well known that it can be diagonalized in $\mathbb{R}$. If $u\in\mathbb{R}^{n}$ is an eigenvector of $B_{y}$ for $y\in\mathcal{Y}$ fixed, with associated eigenvalue $\lambda$ and $\left\lvert u\right\rvert_{2}=1$,
$\lambda=u^{T}B_{y}u=b_{y}\left(\sum_{i=1}^{n}u_{i}\varphi_{i},\sum_{i=1}^{n}u_{i}\varphi_{i}\right)\in\mathbb{R}.$
This equation implies, using the Parseval’s Theorem [Sax02, Theorem 4.2], that
$\lambda\leq\left\lvert\lambda\right\rvert=\left\lvert b_{y}\left(\sum_{i=1}^{n}u_{i}\varphi_{i},\sum_{i=1}^{n}u_{i}\varphi_{i}\right)\right\rvert\leq C_{\operatorname{cont}}\left\lVert\sum_{i=1}^{n}u_{i}\varphi_{i}\right\rVert_{H}^{2}=C_{\operatorname{cont}}\left\lvert u\right\rvert_{2}^{2}=C_{\operatorname{cont}},$
and
$\lambda=b_{y}\left(\sum_{i=1}^{n}u_{i}\varphi_{i},\sum_{i=1}^{n}u_{i}\varphi_{i}\right)\geq C_{\operatorname{coer}}\left\lVert\sum_{i=1}^{n}u_{i}\varphi_{i}\right\rVert_{H}^{2}=C_{\operatorname{coer}}\left\lvert u\right\rvert_{2}^{2}=C_{\operatorname{coer}}.$
Consequently, all the eigenvalues of $B_{y}$ for any $y\in\mathcal{Y}$ are in $[C_{\operatorname{coer}},C_{\operatorname{cont}}]$.
Using the fact that
$\left\lVert\operatorname{I}_{n}-\alpha B_{y}\right\rVert_{2}=\sup\left\{\left\lvert\mu\right\rvert\;\middle|\;\mu\text{ is an eigenvalue of }\operatorname{I}_{n}-\alpha B_{y}\right\},$
we conclude, since $C_{\operatorname{coer}}\leq C_{\operatorname{cont}}$, that for $\alpha\in\left(0,\frac{2}{C_{\operatorname{cont}}}\right)$:
$\left\lVert\operatorname{I}_{n}-\alpha B_{y}\right\rVert_{2}\leq\sup_{\lambda\in[C_{\operatorname{coer}},C_{\operatorname{cont}}]}\left\lvert 1-\alpha\lambda\right\rvert=\max\left\{\left\lvert 1-\alpha C_{\operatorname{cont}}\right\rvert,\left\lvert 1-\alpha C_{\operatorname{coer}}\right\rvert\right\}<1.$
We can approximate the inverse of any $B_{y}$ using the MNN of Theorem 4.6 for $\alpha\in\left(0,\frac{2}{C_{\operatorname{cont}}}\right)$ and $\delta=\max\left\{\left\lvert 1-\alpha C_{\operatorname{cont}}\right\rvert,\left\lvert 1-\alpha C_{\operatorname{coer}}\right\rvert\right\}$.
6 Comparison with previous results for ReLU
activation function
Until now, we have worked with an arbitrary activation function $\varrho$ satisfying the Hypothesis 1. In this section, we focus on the ReLU activation function. The $\operatorname{ReLU}$ activation function, which is given by:
$\operatorname{ReLU}(t)\coloneqq\max\left\{0,t\right\}$
and plotted in Figure 2, satisfies Hypothesis 1, as shown below in Lemma 6.3.
Figure 2: $\operatorname{ReLU}$ plot.
As stated in the introduction, the results are generalizations of the ones in [KPRS22], which focus on $\operatorname{ReLU}$ Neural Networks. Moreover, they are not only more general, but they present more precise bounds on the number of weights and layers for Neural Networks when choosing $\varrho=\operatorname{ReLU}$. We recall that the results obtained in [KPRS22] are the following ones:
Proposition 6.1 ([KPRS22, Proposition 3.7]).
Let $K>0$, $m,n,p\in\mathbb{N}$ and $\varepsilon\in(0,1)$. Then, there exist a $\operatorname{ReLU}$-Neural Network $\Phi^{\operatorname{MULT}_{m,p}^{n}}_{\varepsilon,K}$ and a constant $C>0$ satisfying:
$\left\lVert AB-R\left(\Phi^{\operatorname{MULT}_{m,p}^{n}}_{\varepsilon,K}\right)\left(A^{T}\middle|B\right)\right\rVert_{2}\leq\varepsilon$ (6.1)
for all $A\in\mathbb{R}^{m\times n},B\in\mathbb{R}^{n\times p}$ and satisfying:
$M\left(\Phi^{\operatorname{MULT}_{m,p}^{n}}_{\varepsilon,K}\right)$ $\leq Cmnp\left(\log_{2}\frac{1}{\varepsilon}+\log_{2}\left(mnp\right)+\log_{2}\left\{1,K\right\}\right),$ (6.2)
$L\left(\Phi^{\operatorname{MULT}_{m,p}^{n}}_{\varepsilon,K}\right)$ $\leq C\left(\log_{2}\frac{1}{\varepsilon}+\log_{2}\left(mnp\right)+\log_{2}\max\left\{1,K\right\}\right).$ (6.3)
Theorem 6.2 ([KPRS22, Theorem 3.8]).
Let $\delta\in(0,1)$, $\varepsilon\in\left(0,\frac{1}{4}\right)$ and $n\in\mathbb{N}$. We call:
$m(\varepsilon,\delta)=\frac{\log_{2}(\varepsilon(1-\delta)/2)}{\log_{2}\delta}.$
Then, there exists a $\operatorname{ReLU}$-Matrix Neural Network $\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{*}_{n}}$ and a constant $C>0$ satisfying:
$\left\lVert(A)^{-1}-R\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{*}_{n}}\right)(A)\right\rVert_{2}\leq\varepsilon$ (6.4)
for all $A\in\mathbb{R}^{n\times n}$ such that
$\left\lVert\operatorname{I}_{n}-A\right\rVert_{2}\leq\delta,$
and satisfying:
$M\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{*}_{n}}\right)$ $\leq Cn^{3}m(\varepsilon,\delta)\log_{2}^{2}m(\varepsilon,\delta)\left(\log_{2}\frac{1}{\varepsilon}+\log_{2}m(\varepsilon,\delta)+\log_{2}n\right),$ (6.5)
$L\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{*}_{n}}\right)$ $\leq C\log_{2}m(\varepsilon,\delta)\left(\log_{2}\frac{1}{\varepsilon}+\log_{2}m(\varepsilon,\delta)+\log_{2}n\right).$ (6.6)
Now we analyze the results of our paper in the case of the $\operatorname{ReLU}$ activation function. For that, improving the upper bounds on the number of weights and layers of [EGJS22, Lemma 6.2], we may multiply two numbers using a $\operatorname{ReLU}$ Neural Network:
Lemma 6.3.
Let $K>0$ and $\varrho=\operatorname{ReLU}$. Then, for all $\varepsilon>0$ and for $C=\max\left\{1,K\right\}$ there exist a $\operatorname{ReLU}$ Neural Networks $\Pi_{\varepsilon,K}^{\varrho}$ such that we have the following error estimate:
$\sup\limits_{(x,y)\in[-K,K]^{2}}\left\lvert xy-R(\Pi_{\varepsilon,K}^{\varrho})(x,y)\right\rvert\leq\varepsilon,$ (6.7)
such that for every $x\in\mathbb{R}$ it holds that
$R(\Pi_{\varepsilon,K}^{\varrho})(0,x)=R(\Pi_{\varepsilon,K}^{\varrho})(x,0)=0,$ (6.8)
such that we have the following estimate on the number of weights:
$M\left(\Pi_{\varepsilon,K}^{\varrho}\right)\leq\begin{cases}30\log_{2}K+15\log_{2}\frac{1}{\varepsilon}+25&\text{ if }\varepsilon<\frac{K^{2}}{2},\\ 12&\text{ if }\varepsilon\in\left[\frac{K^{2}}{2},K^{2}\right),\\ 0&\text{ if }\varepsilon\geq K^{2},\end{cases}$ (6.9)
and such that we have the following estimate on the length:
$L\left(\Pi_{\varepsilon,K}^{\varrho}\right)\leq\begin{cases}\log_{2}K+\frac{1}{2}\log_{2}\frac{1}{\varepsilon}+\frac{5}{2}&\text{ if }\varepsilon<\frac{K^{2}}{2},\\ 2&\text{ if }\varepsilon\in\left[\frac{K^{2}}{2},K^{2}\right),\\ 1&\text{ if }\varepsilon\geq K^{2}.\end{cases}$ (6.10)
The proof can be found in Appendix A.
With our constructed MNN and Lemma 6.3 we get:
Corollary 6.4.
Let $K>0$, $m,n,p\in\mathbb{N}$ and $\varepsilon\in(0,1)$. Then, there exists a $\operatorname{ReLU}$-Matrix Neural Network $\Phi^{\operatorname{MULT}_{m,p}^{n}}_{\varepsilon,K}$ and a constant $C>0$ satisfying:
$\left\lVert AB-R\left(\Phi^{\operatorname{MULT}_{m,p}^{n}}_{\varepsilon,K}\right)\left(A^{T}\middle|B\right)\right\rVert_{2}\leq\varepsilon$ (6.11)
for all $A\in\mathbb{R}^{m\times n},B\in\mathbb{R}^{n\times p}$ such that $\left\lVert A\right\rVert_{2},\left\lVert B\right\rVert_{2}\leq K$ and
$M\left(\Phi^{\operatorname{MULT}_{m,p}^{n}}_{\varepsilon,K}\right)$ $\leq C\max\left\{m,n,p\right\}^{\log_{2}7}\left[\log_{2}\frac{1}{\varepsilon}+\log_{2}\left(\max\left\{m,n,p\right\}\right)\right.$ (6.12)
$+\log_{2}\max\left\{1,K\right\}+1\right],$
$L\left(\Phi^{\operatorname{MULT}_{m,p}^{n}}_{\varepsilon,K}\right)$ $\leq C\left[\log_{2}\frac{1}{\varepsilon}+\log_{2}(\max\left\{m,n,p\right\})+\log_{2}\max\left\{1,K\right\}+1\right].$ (6.13)
Proof.
This result is a direct consequence of Proposition 3.2 and Lemma 6.3. In fact, property (6.11) is satisfied if we define:
$\Phi^{\operatorname{MULT}_{m,p}^{n}}_{\varepsilon,K}=\Phi_{\frac{\varepsilon}{\sqrt{mp}},\max\left\{1,K\right\}}^{\operatorname{STR}^{n}_{m,p,\operatorname{ReLU}}}.$
This can be proved by using (3.12) and the fact that:
$\left\lvert A\right\rvert_{\infty}\leq\left\lVert A\right\rVert_{2}\leq\sqrt{mp}\left\lvert A\right\rvert_{\infty}$
for all $A\in\mathbb{R}^{m\times p}$.
We can now use Lemma 6.3 and (3.13) to prove the upper bound on the number of weights: if we call $\mu=\max\left\{m,n,p\right\}$ and $\gamma=\max\left\{1,K\right\}$ for legibility, we get:
$M\left(\Phi^{\operatorname{MULT}_{m,p}^{n}}_{\varepsilon,K}\right)$ $\leq 7\mu^{\log_{2}7}\left(M\left(\Pi_{\frac{\varepsilon}{4\sqrt{mp}\mu^{2}},2\mu\gamma}^{\operatorname{ReLU}}\right)+12\right)-9\mu^{2}$
$\leq 7\mu^{\log_{2}7}\left(30\log_{2}(2\mu\gamma)+15\log_{2}\frac{4\sqrt{mp}\mu^{2}}{\varepsilon}+37\right)-9\mu^{2}$
$\leq C\mu^{\log_{2}7}\left(\log_{2}\frac{1}{\varepsilon}+\log_{2}\mu+\log_{2}\gamma+1\right),$
which implies (6.12). We have used the fact that $\varepsilon<1\leq 8\mu^{4}\sqrt{mp}C^{2}$ to choose the upper bound of the number of weights in Lemma 6.3.
We finish proving (6.13). Setting again $\mu=\max\left\{m,n,p\right\}$ and $\gamma=\max\left\{1,K\right\}$, we extract from Lemma 6.3 and (3.14) that
$L\left(\Phi^{\operatorname{MULT}_{m,p}^{n}}_{\varepsilon,K}\right)$ $\leq L\left(\Pi_{\frac{\varepsilon}{4\mu^{2}},2\mu\gamma}^{\operatorname{ReLU}}\right)+2\left(\log_{2}\mu+2\right)$
$\leq\log_{2}(2\mu\gamma)+\frac{1}{2}\log_{2}\frac{4\mu^{2}}{\varepsilon}+\frac{5}{2}+2\log_{2}\mu+4,$
which implies (6.13). ∎
If we do a one-to-one comparison of Proposition 6.1 and Corollary 6.4, we see that the main differences are that the Strassen method gives a better exponent ($\log_{2}7$) than 3 on the number of weights but with a larger number of layers. This last difference comes from the way we concatenate two Matrix Neural Networks. The concatenation definition given in [KPRS22] for the Strassen Matrix Neural Network sacrifices a better upper bound on the number of weights in order to gain a smaller footprint on the number of layers. In any case, the cost of evaluating a longer Matrix Neural Network is negligible when working with $\operatorname{ReLU}$ as an activation function.
Corollary 6.5.
Let $\delta\in(0,1)$, $\varepsilon\in\left(0,\frac{1}{4}\right)$ and $n\in\mathbb{N}$. We call
$m(\varepsilon,\delta)=\frac{\log_{2}\left(\varepsilon(1-\delta)/2\right)}{\log_{2}\delta}.$
Then, there exists a $\operatorname{ReLU}$-Matrix Neural Network $\Phi_{\varepsilon,\delta}^{\operatorname{INV}_{n}^{*}}$ and a constant $C>0$ satisfying:
$\left\lVert A^{-1}-R\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}_{n}^{*}}\right)(A)\right\rVert_{2}\leq\varepsilon$ (6.14)
for all $A\in\mathbb{R}^{n\times n}$ such that
$\left\lVert\operatorname{I}_{n}-A\right\rVert_{2}\leq\delta,$ (6.15)
and satisfying:
$M\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}_{n}^{*}}\right)$ $\leq Cn^{\log_{2}7}\log_{2}m(\varepsilon,\delta)\left(\log_{2}\frac{1}{\varepsilon}+m(\varepsilon,\delta)+\log_{2}n\right),$ (6.16)
$L\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}_{n}^{*}}\right)$ $\leq C\log_{2}m(\varepsilon,\delta)\left(\log_{2}\frac{1}{\varepsilon}+m(\varepsilon,\delta)+\log_{2}n\right)$ (6.17)
if $m(\varepsilon,\delta)>2$, and satisfying:
$M\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{*}_{n}}\right)$ $=2(n^{2}+n),$ (6.18)
$L\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{*}_{n}}\right)$ $=2$ (6.19)
otherwise.
Note that (6.15) implies that we can use Theorem 4.6 with $\alpha=1$. Also, note that for $\varepsilon\in\left(0,\frac{1}{4}\right)$ and $\delta\in(0,1)$:
$2^{N\left(\frac{\varepsilon}{2},\delta\right)}\leq 2m(\varepsilon,\delta),$ (6.20)
where $N$ is given in (4.20).
Proof.
This result is a direct consequence of Theorem 4.6 and Lemma 6.3.
If we define
$\Phi_{\varepsilon,\delta}^{\operatorname{INV}_{n}^{*}}=\Phi_{\varepsilon,\delta}^{\operatorname{INV}_{n,\operatorname{ReLU}}^{1}},$
we get property (6.14) and also (6.18) and (6.19).
Suppose that $N\left(\frac{\varepsilon}{2},\delta\right)=\left\lceil\log_{2}\max\left\{m(\varepsilon,\delta),2\right\}\right\rceil>1$, that is, $m(\varepsilon,\delta)>2$.
We proceed by giving an upper bound on the number of weights using (4.22):
$\begin{split}M\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}_{n}^{*}}\right)&\leq 14n^{\log_{2}7}\left(N\left(\frac{\varepsilon}{2},\delta\right)-1\right)\left(M\left(\Pi^{\operatorname{ReLU}}_{2^{-2^{N\left(\frac{\varepsilon}{2},\delta\right)}}\frac{\varepsilon}{16n^{3}},2n}\right)+12\right)\\ &+n^{2}\left(L\left(\Pi_{2^{-2^{N\left(\frac{\varepsilon}{2},\delta\right)}}\frac{\varepsilon}{16n^{3}},2n}^{\operatorname{ReLU}}\right)-14N\left(\frac{\varepsilon}{2},\delta\right)+20+2\log_{2}n\right)\\ &+n\left(N\left(\frac{\varepsilon}{2},\delta\right)+1\right).\end{split}$ (6.21)
By (6.9), (6.20) and $\log_{2}\left(\frac{1}{\varepsilon}\right)\geq 1$, we have:
$M\left(\Pi^{\operatorname{ReLU}}_{2^{-2^{N\left(\frac{\varepsilon}{2},\delta\right)}}\frac{\varepsilon}{16n^{3}},2n}\right)$ $\leq 30\log_{2}2n+15\log_{2}\left(2^{2^{N(\frac{\varepsilon}{2},\delta)}}\frac{16n^{3}}{\varepsilon}\right)+25$
$\leq C\left(\log_{2}\frac{1}{\varepsilon}+m(\varepsilon,\delta)+\log_{2}n\right).$
Similarly, by (6.10), (6.20) and $\log_{2}\left(\frac{1}{\varepsilon}\right)\geq 1$, we have:
$L\left(\Pi_{2^{-2^{N\left(\frac{\varepsilon}{2},\delta\right)}}\frac{\varepsilon}{16n^{3}},2n}^{\operatorname{ReLU}}\right)$ $\leq\log_{2}2n+\frac{1}{2}\log_{2}\left(2^{2^{N(\frac{\varepsilon}{2},\delta)}}\frac{16n^{3}}{\varepsilon}\right)+\frac{5}{2}$
$\leq C\left(\log_{2}\frac{1}{\varepsilon}+m(\varepsilon,\delta)+\log_{2}n\right).$
Introducing those bounds into (6.21) we get:
$M\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}_{n}^{*}}\right)$ $\leq C\left(n^{\log_{2}7}N\left(\frac{\varepsilon}{2},\delta\right)+n^{2}\right)\left(\log_{2}\frac{1}{\varepsilon}+m(\varepsilon,\delta)+\log_{2}n\right)$
$+CnN\left(\frac{\varepsilon}{2},\delta\right)$
$\leq Cn^{\log_{2}7}N\left(\frac{\varepsilon}{2},\delta\right)\left(\log_{2}\frac{1}{\varepsilon}+m(\varepsilon,\delta)+\log_{2}n\right),$
which proves (6.16).
Finally, let us prove (6.17). Using (4.23) and (6.10):
$L\left(\Phi^{\operatorname{INV}^{*}_{n}}_{\varepsilon,\delta}\right)$ $\leq N\left(\frac{\varepsilon}{2},\delta\right)\left[2\log_{2}n+5+L\left(\Pi_{2^{-2^{N\left(\frac{\varepsilon}{2},\delta\right)}}\frac{\varepsilon}{16n^{3}},2n}^{\varrho}\right)\right]$
$\leq CN\left(\frac{\varepsilon}{2},\delta\right)\left(\log_{2}\frac{1}{\varepsilon}+m(\varepsilon,\delta)+\log_{2}n\right).$
This result presents various improvements with respect to Theorem 6.2. Although the number of layers is asymptotically worse, the upper bound on the number of weights on (6.16) is an improvement over (6.5). The exponent on the dimension of the input matrices is reduced from 3 to $\log_{2}7$ and the factor is $\log_{2}m(\varepsilon,\delta)$ instead of $m(\varepsilon,\delta)\log_{2}^{2}(m(\varepsilon,\delta))$. In parentheses, even if we have $m(\varepsilon,\delta)$ instead of $\log_{2}m(\varepsilon,\delta)$, it is still an improvement if we compare the expanded expression $m(\varepsilon,\delta)\log_{2}(m(\varepsilon,\delta))$ to $m(\varepsilon,\delta)\log_{2}(m(\varepsilon,\delta))\log_{2}^{2}(m(\varepsilon,\delta))$. The first change comes from the use of the Strassen algorithm for matrix multiplication, while the second is due to the use of the identity (4.2).
7 Conclusions
We have constructed two important Neural Networks approximating the matrix multiplication and inversion operators. We have been able to do so without restricting ourselves to a fixed activation function. Also, we bound their number of weights and layers, and by using the Strassen algorithm we have reduced a potential cubic dependency on the dimension to a $\log_{2}7$ exponent. Rewriting the Neumann partial sum, we reduce the upper bounds on the number of parameters by reducing the number of matrix multiplication apparitions. In Section 6, we show the size improvements of these Neural Networks with respect to previous approximations to the matrix multiplication and inversion operators. We also explain how the Neural Network that approximates matrix inversion operator is useful for solving parametric elliptic PDEs.
As explained, there is a high interest in finding optimal upper bounds on the number of parameters for approximating maps with Neural Networks. A remaining challenge is then finding explicitly the smallest Neural Neural tasked for matrix multiplication or inversion approximation while giving its explicit construction. The answer is unknown even for some simple activation function such as $\operatorname{ReLU}$.
Appendix A Proof of Lemma 6.3
To prove Lemma 6.3, we use the following lemma:
Lemma A.1.
For every $\varepsilon>0$, there is a $\operatorname{ReLU}$ Neural Network $\Phi^{\operatorname{SQR}}_{\varepsilon}$ with input and output dimensions 1 such that we have the following error estimate:
$\sup\limits_{t\in[0,1]}\left\lvert t^{2}-R\left(\Phi^{\operatorname{SQR}}_{\varepsilon}\right)(t)\right\rvert\leq\varepsilon,$ (A.1)
such that we have the following equality:
$R\left(\Phi^{\operatorname{SQR}}_{\varepsilon}\right)(0)=0,$ (A.2)
such that we have the following estimates on the initial and final weights if $\varepsilon<\frac{1}{4}$:
$M_{1}\left(\Phi^{\operatorname{SQR}}_{\varepsilon}\right)=6,\ \ \ M_{L\left(\Phi^{\operatorname{SQR}}_{\varepsilon}\right)}\left(\Phi^{\operatorname{SQR}}_{\varepsilon}\right)=4,$ (A.3)
such that we have the following estimates on the number of weights:
$M\left(\Phi^{\operatorname{SQR}}_{\varepsilon}\right)=\begin{cases}\max\left\{15\left\lceil\frac{1}{2}\log_{2}\frac{1}{\varepsilon}\right\rceil-20,1\right\}&\text{ if }\varepsilon<1,\\ 0&\text{ if }\varepsilon\geq 1,\end{cases}$ (A.4)
and such that we have the following estimate on its length:
$L\left(\Phi^{\operatorname{SQR}}_{\varepsilon}\right)=\max\left\{\left\lceil\frac{1}{2}\log_{2}\frac{1}{\varepsilon}\right\rceil,1\right\}.$ (A.5)
Lemma A.1 is proved in [EGJS22, Lemma 6.1] (a proof that is based on [Yar17, Proposition 3]). Note that for $\varepsilon\geq 1$ we can consider the constant function 00 as a neural network, and for $\varepsilon\in[\frac{1}{4},1)$ we can consider the neural network $(0,0,\operatorname{id})$.
In order to prove Lemma 6.3, instead of using
$xy=\frac{1}{2}(x+y)^{2}-x^{2}-y^{2}$
as in [EGJS22], we use that
$xy=K^{2}\left(\left(\frac{x+y}{2K}\right)^{2}-\left(\frac{x-y}{2K}\right)^{2}\right)$ (A.6)
for all $K>0$. This simplifies the implicit Neural Network as well as the resulting upper bounds on the number of weights and layers. We are now ready to prove Lemma 6.3:
Proof of Lemma 6.3.
Let us fix $K>0$. We divide the proof into two cases: $\varepsilon\geq K^{2}$ and $\varepsilon<K^{2}$.
The case $\varepsilon\geq K^{2}$ follows immediately. In fact, from
$\sup_{x,y\in[-K,K]}\left\lvert xy-0\right\rvert\leq K^{2},$
we define $\Pi^{\operatorname{ReLU}}_{\varepsilon,K}=\left(\left(\begin{pmatrix}0&0\end{pmatrix},0,\operatorname{id}\right)\right)$, which satisfies all the desired properties.
We now suppose that $\varepsilon<K^{2}$. By Lemma A.1 and (A.6), we know that:
$\left\lvert xy-K^{2}\left(R\left(\Phi^{\operatorname{SQR}}_{\varepsilon}\right)\left(\frac{\left\lvert x+y\right\rvert}{2K}\right)-R\left(\Phi^{\operatorname{SQR}}_{\varepsilon}\right)\left(\frac{\left\lvert x-y\right\rvert}{2K}\right)\right)\right\rvert\leq 2\varepsilon K^{2}.$
We now define $((A^{1},b^{1},\alpha^{1}),\dots,(A^{L},b^{L},\alpha^{L}))$ as follows:
$((A^{1},b^{1},\alpha^{1}),\dots,(A^{L},b^{L},\alpha^{L}))\coloneq\mathcal{P}\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}},\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right).$
We introduce the auxiliary Neural Networks $\Phi^{\operatorname{ABS}}$ and $\Phi^{\operatorname{SUM}}$:
$\Phi^{\operatorname{ABS}}$ :-$\coloneq\left(\left(\frac{1}{2K}\begin{pmatrix}1&1\\ -1&-1\\ 1&-1\\ -1&1\end{pmatrix},0,\operatorname{ReLU}\times\operatorname{ReLU}\times\operatorname{ReLU}\times\operatorname{ReLU}\right),\right.:$
[…]
$\Phi^{\operatorname{SUB}}$ :-$\coloneq\left(\begin{pmatrix}K^{2}&-K^{2}\end{pmatrix},0,\operatorname{id}\right).:$
Note that these Neural Networks satisfy:
$R\left(\Phi^{\operatorname{ABS}}\right)\begin{pmatrix}x\\ y\end{pmatrix}$ $=\begin{pmatrix}\operatorname{ReLU}\left(\frac{x+y}{2K}\right)+\operatorname{ReLU}\left(\frac{-x-y}{2K}\right)\\ \operatorname{ReLU}\left(\frac{x-y}{2K}\right)+\operatorname{ReLU}\left(-\frac{x-y}{2K}\right)\end{pmatrix}=\frac{1}{2K}\begin{pmatrix}\left\lvert x+y\right\rvert\\ \left\lvert x-y\right\rvert\end{pmatrix},$
$R\left(\Phi^{\operatorname{SUB}}\right)\begin{pmatrix}x\\ y\end{pmatrix}$ $=K^{2}(x-y).$
Then, the target Neural Network is of the form $\Phi^{\operatorname{SUB}}\circ\mathcal{P}\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}},\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right)\circ\Phi^{\operatorname{ABS}}$, but to keep it more compact we define it as:
[…]
where $\alpha=\operatorname{ReLU}\times\operatorname{ReLU}\times\operatorname{ReLU}\times\operatorname{ReLU}$. If $L=1$, we understand the right hand side of the previous definition as:
[…]
We deduce from (A.1) and (A.2) that (6.7) and (6.8) are satisfied respectively.
Let us now prove the upper bound (6.9) on the number of weights. We have, for $\varepsilon<\frac{K^{2}}{2}$ (which implies $L>1$),
$\begin{split}M\left(\Pi^{\varrho}_{\varepsilon,K}\right)&=\operatorname{n_{W}}\left(\frac{1}{2K}\begin{pmatrix}1&1\\ -1&-1\\ 1&-1\\ -1&1\end{pmatrix}\right)\\ &+\operatorname{n_{W}}\left(A^{1}\begin{pmatrix}1&1&0&0\\ 0&0&1&1\end{pmatrix}\right)+\operatorname{n_{W}}\left(b^{1}\right)\\ &+M\left(\mathcal{P}\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}},\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right)\right)\\ &-M_{1}\left(\mathcal{P}\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}},\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right)\right)-M_{L}\left(\mathcal{P}\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}},\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right)\right)\\ &+\operatorname{n_{W}}\left(\begin{pmatrix}K^{2}&-K^{2}\end{pmatrix}A^{L}\right)+\operatorname{n_{W}}\left(\begin{pmatrix}K^{2}&-K^{2}\end{pmatrix}b^{1}\right).\end{split}$ (A.8)
Let us estimate the terms in the right hand side of (A.8). To begin with,
$\operatorname{n_{W}}\left(\frac{1}{2K}\begin{pmatrix}1&1\\ -1&-1\\ 1&-1\\ -1&1\end{pmatrix}\right)=8.$ (A.9)
To continue with, note that, considering (A.7):
$\operatorname{n_{W}}\left(A^{1}\begin{pmatrix}1&1&0&0\\ 0&0&1&1\end{pmatrix}\right)+\operatorname{n_{W}}\left(b^{1}\right)=4M_{1}\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right).$
Also, by the nature of parallelization:
$M_{1}\left(\mathcal{P}\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}},\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right)\right)=2M_{1}\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right).$
Thus, from the previous two equations and (A.3) we get that:
$\operatorname{n_{W}}\left(A^{1}\begin{pmatrix}1&1&0&0\\ 0&0&1&1\end{pmatrix}\right)+\operatorname{n_{W}}\left(b^{1}\right)-M_{1}\left(\mathcal{P}\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}},\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right)\right)\\ =2M_{1}\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right)=12.$ (A.10)
Similarly, for the last layer, we obtain that:
$\operatorname{n_{W}}\left(\begin{pmatrix}K^{2}&-K^{2}\end{pmatrix}A^{L}\right)+\operatorname{n_{W}}\left(\begin{pmatrix}K^{2}&-K^{2}\end{pmatrix}b^{1}\right)-M_{L}\left(\mathcal{P}\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}},\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right)\right)\\ \leq M_{L}\left(\mathcal{P}\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}},\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right)\right)-M_{L}\left(\mathcal{P}\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}},\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right)\right)=0.$ (A.11)
Finally, let us estimate the main weights. Considering (A.4) we obtain that:
$\begin{split}M\left(\mathcal{P}\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}},\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right)\right)=2M\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right)&=30\left\lceil\frac{1}{2}\log_{2}\frac{2K^{2}}{\varepsilon}\right\rceil-40\\ &\leq 15\log_{2}\frac{2K^{2}}{\varepsilon}-10\\ &=30\log_{2}K+15\log_{2}\frac{1}{\varepsilon}+5,\end{split}$ (A.12)
where we have used $\left\lceil x\right\rceil\leq x+1$ for all $x\in\mathbb{R}$. Thus, we obtain (6.9) from (A.8), (A.9), (A.10), (A.11) and (A.12).
When $\varepsilon\in\left[\frac{K^{2}}{2},K^{2}\right)$ (which implies $L=1$), we get as before
$M\left(\Pi_{\varepsilon,K}^{\varrho}\right)\leq 12.$
This proves (6.9).
Finally, the number of layers follows directly from (A.5). Indeed,
$\begin{split}L\left(\Pi^{\varrho}_{\varepsilon,K}\right)&=L+1=L\left(\Phi^{\operatorname{SQR}}_{\frac{\varepsilon}{2K^{2}}}\right)+1\\ &=\left\lceil\frac{1}{2}\log_{2}\frac{2K^{2}}{\varepsilon}\right\rceil+1\\ &\leq\log_{2}K+\frac{1}{2}\log_{2}\frac{1}{\varepsilon}+\frac{5}{2},\end{split}$
if $\varepsilon<\frac{K^{2}}{2}$, and
$L\left(\Pi^{\varrho}_{\varepsilon,K}\right)=2$
if $\varepsilon\in\left[\frac{K^{2}}{2},K^{2}\right)$. This shows (6.10). ∎